### v1

# A Simple First-Order Algorithm for Full-Rank Equality Constrained Optimization

Abstract
A very simple first-order algorithm is proposed for solving nonlinear optimization problems with nonlinear equality constraints. This algorithm adaptively selects steps in the plane tangent to the constraints or steps that reduce infeasibility, without using a merit function or filter. The tangent steps are based on the AdaGrad method for unconstrained minimization. The objective function is never evaluated by the algorithm, making it suitable for noisy problems. Its worst-case evaluation complexity is analyzed, yielding a global convergence rate in ${\pazocal{O}}(1/\sqrt{k})$, which matches the best known rate of first-order methods for unconstrained problems. Numerical experiments are presented suggesting that the performance of the algorithm is comparable to that of first-order methods for unconstrained problems, and that its reliability is remarkably stable in the presence of noise on the gradient.
Keywords: Equality constrained optimization, objective-function-free optimization (OFFO), first-order methods, AdaGrad, evaluation complexity.
1 Introduction
Solving nonlinear problems with nonlinear equality constraints is a central part of more general constrained optimization, and the purpose of this paper is to contribute to the subject by proposing a very simple algorithm for the case where the constraints’ Jacobian is full-rank. We introduce a method drawing its inspiration both from the “trust-funnel” approaches [17, 8] and the objective-function free (OFFO) first-order algorithms that have become very popular because of their successful use in deep learning applications. Trust-funnel methods themselves exploit the older idea of decomposing the step computed within an optimization algorithm for constrained problems into a “tangential” and a “normal” components. Advocated in particular by Byrd and Omojokun [24] for sequential quadratic programming, this last idea is to consider that a useful minimization step should, on the one hand, improve the value of the objective function without deteriorating the constraints’s violation too much (this is the tangential step, because it can be interpreted as a step (nearly) tangent to the constraint’s surface) and, on the other hand, improve feasibility (the normal step, mostly orthogonal to the constraints’ surface). This technique has been successfully used by many algorithms, and in particular by trust-funnel ones. In this case, the two types of steps are treated separately and a choice between normal and tangential steps is made at every iteration by comparing constraint violation (primal feasibility) with some dual optimality measure. Our new proposed algorithm uses a similar mechanism, although considerably simplified. Our second source of inspiration is the thriving field of first-order methods for unconstrained problems, and, more specifically, the well-known and well understood AdaGrad [10, 29, 9, 19] algorithm. Such methods are attractive because they avoid computing the value of the objective function entirely (they only rely on gradients). This feature makes them quite robust in the presence of noise on the gradient, as is commonly the case when its evaluation involves subsampling. Again, our new method exploits this OFFO [^1] Objective-Function-Free Optimization. strategy to achieve robustness. In effect, it consists of a simplified trust-funnel method in which the tangential step is computed using the AdaGrad algorithm.
Other OFFO methods for constrained optimization have been proposed. The stochastic algorithm proposed in [1] uses a combined tangential and normal step for a quadratic model of a merit function. A (nontrivial) stochastic analysis of its convergence (albeit not its worst-case complexity) is provided. It also allows the use of second-order information, at the cost of a somewhat complicated algorithm. [7] extends this proposal to handle inequality constraints and provides an updated convergence analysis. The algorithm of [11] is more similar to ours, but differs in significant aspects. Although it uses tangential and normal steps, it does so at every iteration. The step in the Jacobian’s nullspace is computed by a standard trust-region technique and a special rule is provided to update its radius. It also relies on an explicit penalty parameter with is updated in the course of the algorithm. Convergence of the algorithm is also analyzed in a stochastic setting. The method of [4] is inspired by filter methods (another class of methods for nonlinear optimization [13, 14, 12]) and also alternates between steps which attempts to reduce constraints’ infeasibilities and “optimality” steps improving the objective function value, albeit ignoring the presence of constraints in the second case. Good numerical results are reported on deep learning problems, but no convergence theory or complexity analysis is provided [^2] Since the optimality steps ignore the constraints’ geometry, it seems possible that their effect might significantly decrease feasibility and that cycling could occur..
In view of these comments, we summarize our contributions as follows.
  1. 1.
    In order to solve problem (2.1), we propose an “adaptive’ switching” algorithm, dubbed ADSWITCH, which can be viewed as a simplified trust-funnel method. ADSWITCH uses the purely OFFO first-order AdaGrad algorithm in the plane tangent to the constraints.
  2. 2.
    We provide a complete worst-case complexity analysis for the case where the objective’s gradient and the constraints are deterministic, under assumptions similar to those used in [1, 7].
  3. 3.
    We illustrate the practical behaviour of the method on a subset of the CUTEst problems [15] as provided by the S2MPJ environment [21].
The paper is organized as follows. The ADSWITCH algorithm is introduced in Section 2 together with the necessary concepts and notation. Section 3 contains the relevant complexity analysis. Numerical illustration is finally proposed in Section 4 while some conclusions and perspectives are discussed in Section 5.
Notations: In what follows, $\|\cdot\|$ denotes the Euclidean norm on the relevant space, $\sigma_{\min}(M)$ is the smallest singular value of the matrix $M$.
2 An adaptative switching algorithm
We propose solving the smooth equality constrained problem
$\min_{x\in\hbox{\footnotesize I\thinspace R}^{n}}f(x)\qquad\text{such that}\qquad c(x)=0,$ (2.1)
using the purely first-order ADSWITCH method, which we now describe after establishing some notation. In (2.1), $f$ is a smooth function from $\hbox{I\thinspace R}^{n}$ into IR and $c$ is a smooth function from $\hbox{I\thinspace R}^{n}$ into $\hbox{I\thinspace R}^{m}$ ($m\leq n$). For a given vector $x$, we assume that we can compute the gradient $g(x)=\nabla_{x}f(x)$, as well as the (exact) Jacobian of the constraints at $x$
$J(x)=\nabla c(x)\in\hbox{I\thinspace R}^{m\times n}.$
Assuming $J(x)$ has full rank, we may also compute the orthogonal projection onto its nullspace by
$P_{T}(x)\stackrel{{\scriptstyle\rm def}}{{=}}I-J(x)^{T}\big(J(x)J(x)^{T}\big)^{-1}J(x).$
Given this projection, the projected gradient is then
$g_{T}(x)\stackrel{{\scriptstyle\rm def}}{{=}}P_{T}(x)\,g(x)$ (2.2)
We may also define the (exact) least-squares Lagrange multiplier $\widehat{\lambda}(x)$ by
$\big(J(x)J(x)^{T}\big)\,\widehat{\lambda}(x)\ =\ -\,J(x)\,\nabla f(x),$ (2.3)
the standard Lagrangian
$L(x,\lambda)=f(x)+\lambda^{T}c(x)$ (2.4)
and, for $\rho$ fixed, the associated “augmented-Lagrangian-like” Lyapunov function
$\psi_{\rho}(x,\lambda)\stackrel{{\scriptstyle\rm def}}{{=}}L(x,\lambda)+\rho\|c(x)\|,$ (2.5)
which we often abbreviate for $\lambda=\widehat{\lambda}(x)$, as
$\psi(x)\stackrel{{\scriptstyle\rm def}}{{=}}\psi_{\rho}\big(x,\widehat{\lambda}(x)\big).$ (2.6)
The definitions imply the important properties that
$g_{T}(x)=\nabla f(x)+J(x)^{T}\widehat{\lambda}(x)\;\;\mbox{ and }\;\;J(x)\,g_{T}(x)=0,$ (2.7)
and also that, for $\lambda$ fixed to $\widehat{\lambda}(x)$,
$\nabla_{x}L\big(x,\lambda\big)=g_{T}(x).$ (2.8)
By construction, we also have that
$c(x)^{T}J(x)g_{T}(x)=0.$ (2.9)
We are now in position to specify ADSWITCH in detail 2. As announced in the introduction, the algorithm uses two types of steps: “tangential steps” attempt to reduce the objective-function value (without evaluating it!) in the nullspace of the constraints, while “normal steps” aim at reducing infeasibility. Which type of step is used at a given iteration depends on the respective values of the tangential step and constraint violation.
Algorithm 2.1: ADSWITCH
Step 0: Initialization. A starting point $x_{0}$ is given, together with constants $\omega>0$, $\theta>1$, $\beta,\eta,\varsigma\in(0,1]$. Set $\Gamma_{0}=0$ and $k=0$. Step 1: Evaluations. Evaluate $c_{k}=c(x_{k})$, $J_{k}=J(x_{k})$ and $g_{k}=\nabla_{x}f(x_{k})$.
Set $g_{T,k}=P_{T}(x_{k})g_{k}$,
$\Gamma_{k}^{+}=\Gamma_{k}+\|g_{T,k}\|^{2},$ (2.10) and $\alpha_{T,k}=\frac{\eta}{\sqrt{\Gamma_{k}^{+}+\varsigma}}$ (2.11)
Step 2: Tangential step: If $\|c_{k}\|\leq\beta\alpha_{T,k}\|g_{T,k}\|$ (2.12) then set $x_{k+1}=x_{k}-\alpha_{T,k}g_{T,k}$ (2.13) and $\Gamma_{k+1}=\Gamma_{k}^{+},$ (2.14) and go to Step 4. Step 3: Normal step: Otherwise, find a step $s_{N,k}$ in the range space of $J_{k}^{T}$ such that $\|s_{N,k}\|\leq\theta\|c_{k}\|$ (2.15) and there exists $\kappa_{\mbox{\tiny nrm}}>0$ independent of $k$ such that ${\scriptstyle\frac{1}{2}}\|c(x_{k}+s_{N,k})\|^{2}\leq{\scriptstyle\frac{1}{2}}\|c_{k}\|^{2}-\kappa_{\mbox{\tiny nrm}}\|J_{k}^{T}c_{k}\|^{2},$ (2.16) and set $x_{k+1}=x_{k}+s_{N,k}.$ (2.17) Step 4: Loop. Increment $k$ by one and go to Step 2.
  • The tangential step is the standard AdaGrad-norm [29] step in the nullspace of the Jacobian.
  • The statement of the procedure to find the normal step in Step 3 may seem abstract because $\kappa_{\mbox{\tiny nrm}}$ may not be known, but, fortunately, there exist several standard techniques to achieve (2.15)-(2.16).
    A first technique is to use a standard Armijo backtracking linesearch along the steepest-descent direction $J_{k}^{T}c_{k}$ (see Lemma 3.3). It is well-known that (2.16) holds in this case with $\kappa_{\mbox{\tiny nrm}}=1/L_{JTc}$ (see [3, Lemma 2.2.1], for instance). More generally, any “gradient related” direction $d_{k}$, that is any $d_{k}$ satisfying (2.15) and $d_{k}^{T}J_{k}^{T}c_{k}\leq-\kappa_{\mbox{\tiny grl}}\|J_{k}^{T}c_{k}\|^{2}$ for some $\kappa_{\mbox{\tiny grl}}>0$, can be used to initiate a backtracking linesearch.
    A second possibility follows along this line and is to select a positive definite matrix $B_{k}$ with bounded condition number, to define $d_{k}=-B_{k}^{-1}J_{k}^{T}c_{k}$ and then to perform a backtracking linesearch along $d_{k}$ (see [5, Section 10.3.1] for a proof that $d_{k}$ is a gradient-related direction), in which case a similar result holds. For instance, using $B_{k}=J_{k}^{T}J_{k}+\delta I_{n}$ for some $\delta\geq 0$ yields the regularized Gauss-Newton step
    $s_{Nk}=-\gamma_{k}(J_{k}^{T}J_{k}+\delta I_{n})^{-1}J_{k}^{T}c_{k}=-\gamma_{k}J_{k}^{T}(J_{k}J_{k}^{T}+\delta I_{m})^{-1}c_{k}$ (2.18)
    where $\gamma_{k}>0$ is chosen by backtracking as large as possible to ensure (2.15) and (2.16). (Note that, for $\delta>0$, $J_{k}^{T}J_{k}+\delta I_{n}$ has bounded condition number when $J_{k}$ is bounded. Moreover this is also the case for $\delta=0$ when $J_{k}$ is full-rank.)
    Suitable variants of trust-region steps are also possible [5, Section 10.3.2], themselves involving several subvariants such as the truncated conjugate gradient [26, 27], CGLS [22] or LSQR [25]. Because iterates generated by these methods are linear combinations of the gradients of ${\scriptstyle\frac{1}{2}}\|c_{k}+J_{k}s\|^{2}$, this guarantees that the final step belongs to the range of $J_{k}^{T}$.
  • We have mentioned an explicit formula for the projection $P_{T}(x)$, but it can also be computed using a rank-revealing QR factorization of $J_{k}^{T}$, the last columns of $Q$ then providing a basis for the nullspace of $J_{k}$. In this case, if $Q_{k}$ and $R_{k}$ are the computed factors, then
    $g_{T,k}=Q_{k,[1:n,m+1:n]}Q_{k,[1:n,m+1:n]}^{T}g_{k}$
    and the Newton step (2.18) is given by
    $s_{N,k}=\left\{\begin{array}[]{ll}-Q_{k,[1:n,1:r]}R_{k,[1:r,1:r]}^{-T}c_{k}&\;\;\mbox{if }\;\;\delta=0,\\ -Q_{k,[1:n,1:m]}R_{k,[1:m,1:m]}(R_{k,[1:m,1:m]}^{T}R_{k,[1:m,1:m]}+\delta I_{m})^{-1}c_{k}&\;\;\mbox{otherwise.}\end{array}\right$. (2.19)
  • Observe that the ADSWITCH algorithm does not involve any explicit merit function or filter to control overall progress, but rely on the simple switching condition (2.12) (hence its name). As we will see below, an implicit merit function (a modified augmented-Lagrangian with suitably chosen parameters) is however crucial for our theoretical argument.
  • ADSWITCH differs in several aspects from the trust-funnel methods of [17, 8]. The first is that it avoids evaluating the objective function in order to improve its reliability in the presence of noise. The second is that it also avoids estimating Lagrange multipliers, but at the cost of computing the exact projection $P_{T}(\cdot)$. The third is that the trust-region approach, which is explicit in trust-funnel methods, is only implicit in ADSWITCH. Indeed it has been argued in [18] that the (unconstrained) AdaGrad algorithm can be viewed as a trust-region method with a specific radius management. Moreover, (2.15) can also be viewed as a trust-region constraint (such a condition is also imposed in [17]), but without the need to explicitly update its radius. Finally, ADSWITCH [^3] In its present incarnation. is more limited than trust-funnel methods in that it does not handle approximate projections in the tangent plane.
In what follows, we denote by $\{k_{\tau}\}\subseteq\{k\}$ the index subsequence of “tangential iterations”, that is iterations where (2.12) holds. We also denote $\{k_{\nu}\}=\{k\}\setminus\{k_{\tau}\}$ the index subsequence of “normal iterations”.
3 Complexity analysis
3.1 Assumptions
Our analysis uses the following assumptions.
AS.0:
$f$ and $c$ are continuously differentiable on $\hbox{I\thinspace R}^{n}$.
AS.1:
For all $x\in\hbox{I\thinspace R}^{n}$, $f(x)\geq f_{\rm low}$.
AS.2:
For all $x\in\hbox{I\thinspace R}^{n}$, $\|g(x)\|\leq\kappa_{g}$ where $\kappa_{g}\geq\eta\beta$.
AS.3:
For all $x\in\hbox{I\thinspace R}^{n}$, $\|c(x)\|\leq\kappa_{c}$.
AS.4:
For all $x\in\hbox{I\thinspace R}^{n}$, $\|J(x)\|\leq\kappa_{J}$
AS.5:
For all $x\in\hbox{I\thinspace R}^{n}$, $\sigma_{\min}(J(x))\geq\sigma_{0}\in(0,1]$.
AS.6:
The gradient $g(x)$ is globally Lipschitz continuous (with constant $L_{g}$).
AS.7:
The Jacobian $J(x)$ is globally Lipschitz continuous (with constant $L_{J}$).
Observe that AS.1, AS.2, AS.3 and AS.4 are automatically satisfied if the iterates $x_{k}$ remain in a bounded domain of $\hbox{I\thinspace R}^{n}$. Also note that
  • AS.5 ensures that $P_{T}(x_{k})$, $g_{T,k}$ and $\widehat{\lambda}(x)$ are well-defined.
  • AS.3, AS.6 and AS.7 ensure that $\widehat{\lambda}(x)$ is Lipschitz continuous (with constant $L_{\lambda}$).
  • AS.4 ensures that $c$ is Lipschitz continuous (with constant $L_{c}$).
[…]
Note that these assumptions are very similar to those of [1, 7]. In particular, AS.5 is also part of Assumption 1 in this last reference.
We finally assume, without loss of generality [^4] The parameter $\varsigma$ may be re-adjusted downwards at the first tangential iteration, if necessary., that
$\Gamma_{k_{\tau}}\geq\varsigma\;\;\;\;(\tau\geq 0).$ (3.1)
3.2 Tangential steps
Our analysis hinges on the fact that first-order descent can be shown on the Lyapunov $\psi(x)$, both for tangential and normal steps, despite the fact that neither $\widehat{\lambda}(x_{k})$ or $\rho$ (which we still need to define) appears in the algorithm. We start by considering tangential steps.
Lemma 3.1
Suppose that AS.4–AS.7 hold. Then $\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})\leq-\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|^{2}+\eta\,\kappa_{\mbox{\tiny tan}}\,\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$ (3.2) where $\kappa_{\mbox{\tiny tan}}=\frac{1}{2}\Big(L_{L}+\rho L_{c}+\beta L_{\lambda}+L_{\lambda}L_{c}\Big).$
  • Proof.    We have that
    $\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})$ $=\underbrace{\psi_{\rho}(x_{k_{\tau}+1},\widehat{\lambda}(x_{k_{\tau}}))-\psi_{\rho}(x_{k_{\tau}},\widehat{\lambda}(x_{k_{\tau}}))}_{\Delta_{x}}$ (3.3)
    $\;\;\;\;\;\;\;\;\;\;\;\;+\underbrace{\psi_{\rho}(x_{k_{\tau}+1},\widehat{\lambda}(x_{k_{\tau}+1}))-\psi_{\rho}(x_{k_{\tau}+1},\widehat{\lambda}(x_{k_{\tau}}))}_{\Delta_{\lambda}}.$
[…]
    $\Delta_{x}=-\nabla_{x}\psi_{\rho}(x_{k_{\tau}},\widehat{\lambda}(x_{k_{\tau}}))^{T}(\alpha_{T,k_{\tau}}g_{T,k_{\tau}})+r_{0}+\rho(\|c(x_{k_{\tau}+1})\|-\|c(x_{k_{\tau}})\|),$
    with
    $|r_{0}|\leq\frac{L_{L}}{2}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$
    Now
    $\|c(x_{k_{\tau}+1})\|=\|c(x_{k_{\tau}})-\alpha_{T,k_{\tau}}J_{k}g_{T,k_{\tau}}+r_{1}\|\leq\|c(x_{k_{\tau}})\|+\|r_{1}\|$
    with
    $\|r_{1}\|\leq\frac{L_{c}}{2}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2},$
    giving
    $\|c(x_{k_{\tau}+1})\|-\|c(x_{k_{\tau}})\|\leq\frac{L_{c}}{2}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$ (3.4)
    Successively using (2.8) and (2.9), we obtain that
    $\Delta_{x}$ $=-\alpha_{T,k_{\tau}}g_{T,k_{\tau}}^{T}g_{T,k_{\tau}}+r_{0}+\frac{\rho L_{c}}{2}\,\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}$ (3.5)
    $\leq-\alpha_{T,k_{\tau}}g_{T,k_{\tau}}^{T}g_{T,k_{\tau}}+\frac{1}{2}\left(L_{L}+\rho L_{c}\right)\,\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}$
    $\leq-\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|^{2}+\frac{1}{2}\left(L_{L}+\rho L_{c}\right)\,\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$
    Now, we may use the Lipschitz continuity of $\widehat{\lambda}$ and $c$ to deduce that
    $\Delta_{\lambda}$ $=c_{k_{\tau}+1}^{T}\big(\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\big)$ (3.6)
    $=(c_{k_{\tau}+1}-c_{k_{\tau}})^{T}\big(\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\big)+c_{k_{\tau}}^{T}\big(\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\big)$
    $\leq\|c_{k_{\tau}}\|\,\|\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\|+\|c_{k_{\tau}+1}-c_{k_{\tau}}\|\,\|\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\|$
    $\leq L_{\lambda}\|c_{k_{\tau}}\|\,\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|+L_{\lambda}L_{c}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2},$
    Taking (2.12) into account gives that $\|c_{k_{\tau}}\|\leq\beta\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|$, we obtain that
    $\Delta_{\lambda}\leq\left(\beta L_{\lambda}+L_{\lambda}L_{c}\right)\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$
    Thus, summing $\Delta_{x}$ and $\Delta_{\lambda}$,
    $\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})\leq$ $-\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|^{2}+\frac{1}{2}(L_{L}+\rho L_{c})\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}$
    $+\left(\beta L_{\lambda}+L_{\lambda}L_{c}\right)\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}$
    and (3.2) follows. $\Box$
The bound (3.2) quantifies the effect of tangential steps on the Lyapunov function, and its right-hand side involves a first-order (descent) term and a second-order perturbation term. We now derive crucial bounds on these terms.
Lemma 3.2
Suppose that AS.2 and AS.5 hold. If we denote $\Gamma_{k_{\tau+1}}=\Gamma_{k_{\tau}}+\|g_{T,k_{\tau}}\|^{2}\;\;\mbox{ and }\;\;\alpha_{T,k_{\tau}}=\frac{\eta}{\sqrt{\Gamma_{k_{\tau+1}}+\varsigma}},$ then, for all $0\leq\tau_{0}\leq\tau_{1}$, $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}\,\|g_{T,k_{\tau}}\|^{2}>\frac{\eta}{2\sqrt{2}}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}$ (3.7) $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}^{2}\,\|g_{T,k_{\tau}}\|^{2}\leq\eta^{2}\,\log\left(\frac{\Gamma_{k_{\tau_{1}+1}}+\varsigma}{\Gamma_{k_{\tau_{0}}}+\varsigma}\right).$ (3.8)
  • Proof.    Let $w_{k_{\tau}+1}=\sqrt{\Gamma_{k_{\tau+1}}+\varsigma}$. The definition (2.11) implies that
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}\,\|g_{T,k_{\tau}}\|^{2}$ $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{\|g_{T,k_{\tau}}\|^{2}}{\sqrt{\Gamma_{k_{\tau+1}}+\varsigma}}$
    $>\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{\|g_{T,k_{\tau}}\|^{2}}{w_{k_{\tau+1}}+w_{k_{\tau}}}$
    $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{w_{k_{\tau+1}}^{2}-w_{k_{\tau}}^{2}}{w_{k_{\tau+1}}+w_{k_{\tau}}}$
    $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}(w_{k_{\tau+1}}-w_{k_{\tau}}),$
    $=\eta(w_{k_{\tau_{1}+1}}-w_{k_{\tau_{0}}}),$
    Now observe that, using (3.1),
    $w_{k_{\tau_{1}+1}}-w_{k_{\tau_{0}}}$ $=\sqrt{\Gamma_{k_{\tau_{1}+1}}+\varsigma}-\sqrt{\Gamma_{k_{\tau_{0}}}+\varsigma}$
    $=\frac{\Gamma_{k_{\tau_{1}+1}}-\Gamma_{k_{\tau_{0}}}}{\sqrt{\Gamma_{k_{\tau_{1}+1}}+\varsigma}+\sqrt{\Gamma_{k_{\tau_{0}}}+\varsigma}}$
    $\geq\frac{\Gamma_{k_{\tau_{1}+1}}-\Gamma_{k_{\tau_{0}}}}{\sqrt{2\Gamma_{k_{\tau_{1}+1}}}+\sqrt{2\Gamma_{k_{\tau_{0}}}}}$
    $>\frac{\Gamma_{k_{\tau_{1}+1}}}{2\sqrt{2\Gamma_{k_{\tau_{1}+1}}}}=\frac{\sqrt{\Gamma_{k_{\tau_{1}+1}}}}{2\sqrt{2}}$
    which gives (3.7). Using the concavity and the increasing nature of the logarithm, we also have from (2.11) that
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}=\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\frac{\|g_{T,k_{\tau}}\|^{2}}{\Gamma_{k_{\tau+1}}+\varsigma}=\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\frac{\Gamma_{k_{\tau+1}}-\Gamma_{k_{\tau}}}{\Gamma_{k_{\tau+1}}+\varsigma}\leq\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\log(\Gamma_{k_{\tau+1}}+\varsigma)-\log(\Gamma_{k_{\tau}}+\varsigma),$
    giving (3.8). $\Box$
3.3 Normal steps
Having considered tangential steps, we now show that a decrease in the Lyapunov function $\psi$ may also be proved if a normal step is taken, provided the penalty parameter $\rho$ is chosen large enough. We start with a fairly simple observation.
Lemma 3.3
Suppose that AS.5 holds and that a normal step is used at iteration $k_{\nu}$. Then $\|c_{k_{\nu}+1}\|\leq\left(1-\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\right)\|c_{k_{\nu}}\|.$ (3.9)
  • Proof.    We have from (2.16) that $\|c_{k_{\nu}+1}\|<\|c_{k_{\nu}}\|$ Then
    $\|c_{k_{\nu}}\|(\|c_{k_{\nu}+1}\|-\|c_{k_{\nu}}\|)\leq(\|c_{k_{\nu}+1}\|+\|c_{k_{\nu}}\|)(\|c_{k_{\nu}+1}\|-\|c_{k_{\nu}}\|)=\|c_{k_{\nu}+1}\|^{2}-\|c_{k_{\nu}}\|^{2}$
    and therefore, using (2.16) again, that
    $\|c_{k_{\nu}+1}\|-\|c_{k_{\nu}}\|\leq-\frac{\kappa_{\mbox{\tiny nrm}}\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|^{2}}{\|c_{k_{\nu}}\|}\leq-\frac{\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|^{2}}{\|c_{k_{\nu}}\|}=-\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|.$
    $\Box$
Lemma 3.4
Suppose that AS.3–AS.7 hold and that a normal step is used at iteration $k_{\nu}$. Define $\rho=\frac{1}{\kappa_{\rm nrm}\,\sigma_{0}^{2}}\Big[\;\kappa_{c}\,\theta\,\,L_{\lambda}\;+\;\frac{1}{2}\,\kappa_{c}\,\theta^{2}\,(L_{L}+L_{\lambda}L_{c})\;+\;\eta\;\Big]$ (3.10) Then we have that $\psi(x_{k_{\nu}+1})-\psi(x_{k_{\nu}})\leq-\eta\|c_{k_{\nu}}\|.$ (3.11)
[…]
    $\Delta_{x}$ $=\psi_{\rho}(x_{k_{\nu}+1},\widehat{\lambda}_{k_{\nu}})-\psi_{\rho}(x_{k_{\nu}},\widehat{\lambda}_{k_{\nu}})$ (3.12)
    $=L(x_{k_{\nu}+1},\widehat{\lambda}_{k_{\nu}})-L(x_{k_{\nu}},\widehat{\lambda}_{k_{\nu}})+\rho\big(\|c_{k_{\nu}+1}\|-\|c_{k_{\nu}}\|\big)$
    $\leq(\nabla_{x}L(x_{k_{\nu}},\widehat{\lambda}_{k_{\nu}})^{T}s_{N,k_{\nu}}+r_{3}-\rho\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|$
    $\leq-\rho\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|+\frac{\theta^{2}L_{L}}{2}\|c_{k_{\nu}}\|^{2}.$
    Using AS.6, the Lipschitz continuity of $\widehat{\lambda}$ and $c$ and AS.5, we then deduce that
    $\Delta_{\lambda}$ $=\psi_{\rho}(x_{k_{\nu}+1},\widehat{\lambda}_{k_{\nu}+1})-\psi_{\rho}(x_{k_{\nu}+1},\widehat{\lambda}_{k_{\nu}})$ (3.13)
    $\leq(\|c_{k_{\nu}}\|+\|c_{k_{\nu}+1}-c_{k_{\nu}}\|)\,\|\widehat{\lambda}_{k_{\nu}+1}-\widehat{\lambda}_{k_{\nu}}\|$
    $\leq L_{\lambda}\,\|s_{N,k_{\nu}}\|\,\|c_{k_{\nu}}\|+\frac{L_{\lambda}L_{c}}{2}\|s_{N,k_{\nu}}\|^{2}$
    $\leq L_{\lambda}\,\theta\|c_{k_{\nu}}\|^{2}+\frac{L_{\lambda}L_{c}\theta^{2}}{2}\|c_{k_{\nu}}\|^{2}$
    and thus, summing (3.12) and (3.13), that
    $\psi(x_{k_{\nu}+1})-\psi(x_{k_{\nu}})\leq-\rho\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|+L_{\lambda}\,\theta\kappa_{c}\,\|c_{k_{\nu}}\|+\frac{1}{2}\left(L_{L}+L_{\lambda}L_{c}\right)\theta^{2}\kappa_{c}\|c_{k_{\nu}}\|.$
    The bound (3.11) then follows from (3.10).
    $\Box$
3.4 Telescoping sum
We now combine the effects of tangential and normal steps to accumulate in a telescoping sum the relevant decreases on $\psi(x)$ across all iterations.
Lemma 3.5
Suppose that AS.0–AS.7 hold. Then $\sqrt{\Gamma_{k_{\tau_{1}+1}}}+\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right),$ (3.14) where $\kappa_{\mbox{\tiny gap}}=\frac{2\sqrt{2}}{\eta}\left(\psi(x_{0})+\frac{\kappa_{J}\kappa_{g}\kappa_{c}}{\sigma_{0}^{2}}+\rho\kappa_{c}-f_{\rm low}\right).$
  • Proof.    Combining (3.2), (3.7) and (3.8), we obtain that
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\left(\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})\right)$ $\leq-\frac{\eta}{2\sqrt{2}}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}+\eta^{2}\kappa_{\mbox{\tiny tan}}\log\left(\frac{\Gamma_{k_{\tau_{1}+1}}+\varsigma}{\Gamma_{k_{\tau_{0}}}+\varsigma}\right)$ (3.15)
    Then, if $\min[k_{\nu_{0}},k_{\tau_{0}}]=0$ and $\max[k_{\nu_{1}},k_{\tau_{1}}]=k$, we have that $w_{k_{\tau_{0}}}=\varsigma$ and $\Gamma_{k_{\tau_{0}}}=0$. The bounds (3.11) and (3.15) therefore give that
    $\psi(x_{k+1})$ $-\psi(x_{0})$ (3.16)
    $=\sum_{\tau=\tau_{0}}^{\tau_{1}}\big(\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})\big)+\sum_{\nu=\nu_{0}}^{\nu_{1}}\big(\psi(x_{k_{\nu}+1})-\psi(x_{k_{\nu}})\big)$
    $\leq-\frac{\eta}{2\sqrt{2}}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}-\eta\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|+\eta^{2}\kappa_{\mbox{\tiny tan}}\log\left(\frac{\Gamma_{k_{\tau_{1}+1}}+\varsigma}{\Gamma_{k_{\tau_{0}}}+\varsigma}\right)$
    $\leq-\frac{\eta}{2\sqrt{2}}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}-\frac{\eta}{2\sqrt{2}}\,\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|+\eta\kappa_{\mbox{\tiny tan}}\log\left(1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right)$
    where we used the fact that $\eta\leq 1$. But (2.3), AS.2, AS.4, and AS.5 ensure that
    $\|\widehat{\lambda}(x)\|\leq\frac{\kappa_{J}}{\sigma_{0}^{2}}\|g(x)\|\leq\frac{\kappa_{J}\kappa_{g}}{\sigma_{0}^{2}}.$
    Hence, using (2.5), AS.1, AS.2, AS.3, we have that
    $\psi(x_{k+1})-\psi(x_{0})\geq f_{\rm low}-\frac{\kappa_{J}\kappa_{g}\kappa_{c}}{\sigma_{0}^{2}}-\rho\kappa_{c}-\psi(x_{0})\stackrel{{\scriptstyle\rm def}}{{=}}-\eta\kappa_{\mbox{\tiny gap}},$
    so that (3.16) implies (3.14). $\Box$
This result is central in our analysis as it provides consistent upper bounds on a quantity related to the optimality mesure $\|g_{T,k}\|+\|c_{k}\|$, which we now exploit separately.
3.5 Tangential complexity
Our next step is to use (3.14) to derive a global rate of convergence over tangential steps. We start by proving a useful technical result.
Lemma 3.6
Suppose that $at\leq b+c\,\log(t)$ for $t\geq 1$. Then $t\leq\max\left[e^{b/c},\frac{4c^{2}}{a^{2}}\right].$
  • Proof.    Suppose that $t>e^{b/c}$. Then $b\leq c\,\log(t)$ and thus (see [28, eq. (14)])
    $at\leq 2c\,\log(t)\leq 2c\,\log(1+t)\leq\frac{2ct}{\sqrt{1+t}}.$
    Hence $a\sqrt{1+t}\leq 2c$, which is to say that $t\leq\left(2c/a\right)^{2}-1$, yielding the desired bound. $\Box$
We now consider the rate of convergence for tangential steps proper.
Lemma 3.7
Suppose that AS.0–AS.7 hold. Then $\sqrt{\Gamma_{k_{\tau_{1}+1}}}\leq\kappa_{T}\stackrel{{\scriptstyle\rm def}}{{=}}\sqrt{\frac{\varsigma}{2}}\max\left[e^{\frac{\kappa_{\mbox{\tiny gap}}}{2\kappa_{\mbox{\tiny tan}}}},\frac{32\kappa_{\mbox{\tiny tan}}^{2}}{\varsigma}\right]$ (3.17) and $\sum_{\tau=\tau_{0}}^{\tau_{1}}\|g_{T,k_{\tau}}\|+\|c_{k_{\tau}}\|\leq\kappa_{T}\sqrt{\tau_{1}+1}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right).$ (3.18)
  • Proof.    The bound (3.14) implies that
    $\sqrt{\Gamma_{k_{\tau_{1}+1}}}\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right).$
    Now (3.1) implies that
    $1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\leq\frac{2}{\varsigma}\Gamma_{k_{\tau_{1}+1}}$
    and thus that
    $\sqrt{\frac{\varsigma}{2}}\sqrt{\frac{2\Gamma_{k_{\tau_{1}+1}}}{\varsigma}}\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(\frac{2\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right)=\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\sqrt{\frac{2\Gamma_{k_{\tau_{1}+1}}}{\varsigma}}\right).$
    Using Lemma 3.6, we then obtain that
    $\sqrt{\Gamma_{k_{\tau_{1}+1}}}=\sqrt{\frac{\varsigma}{2}}\sqrt{\frac{2\Gamma_{k_{\tau_{1}+1}}}{\varsigma}}\leq\sqrt{\frac{\varsigma}{2}}\max\left[e^{\frac{\kappa_{\mbox{\tiny gap}}}{2\kappa_{\mbox{\tiny tan}}}},\frac{32\kappa_{\mbox{\tiny tan}}^{2}}{\varsigma}\right]$
    This is (3.17). We may now invoke the inequality
    $\sum_{j=0}^{k}a_{j}\leq\sqrt{k+1}\sqrt{\sum_{j=0}^{k}a_{j}^{2}}$
    for nonnegative $\{a_{j}\}_{j=0}^{k}$ to deduce from (2.10) and (3.17) that
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\|g_{T,k_{\tau}}\|\leq\sqrt{\tau_{1}+1}\,\sqrt{\sum_{\tau=\tau_{0}}^{\tau_{1}}\|g_{T,k_{\tau}}\|^{2}}=\sqrt{\tau_{1}+1}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}\leq\sqrt{\tau_{1}+1}\,\kappa_{T}.$ (3.19)
    Using the switching condition (2.12) and the fact that $\alpha_{T,k_{\tau}}\leq\eta/\sqrt{\varsigma}$, we also deduce that
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\|c_{k_{\tau}}\|\leq\sum_{\tau=\tau_{0}}^{\tau_{1}}\beta\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|\leq\frac{\beta\eta}{\sqrt{\varsigma}}\sum_{\tau=\tau_{0}}^{\tau_{1}}\|g_{T,k_{\tau}}\|\leq\frac{\beta\eta\kappa_{T}\sqrt{\tau_{1}+1}}{\sqrt{\varsigma}}.$
    Summing this bound with (3.19) then gives (3.18). $\Box$
Note that a marginally tighter bound on $\Gamma_{k_{\tau_{1}}}$ can obtained, as in [2], by using the Lambert $W_{-1}$ function (see [6]) instead of applying Lemma 3.6, but we have chosen the latter for simplicity of exposition.
3.6 Normal complexity
The analysis of the complexity of normal step also uses the switching condition, but in the other direction.
Lemma 3.8
Suppose that AS.0–AS.7 hold. Then $\sum_{\nu=\nu_{0}}^{\nu_{1}}\|g_{T,k_{\nu}}\|+\|c_{k_{\nu}}\|<\kappa_{N}\,\left(1+\frac{2\sqrt{2}}{\beta\eta}\right),$ (3.20) where $\kappa_{N}=\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(\frac{2\kappa_{T}}{\varsigma}\right).$ (3.21)
  • Proof.    The bound (3.14) ensures that
    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right)$ (3.22)
    where $k_{\tau_{1}}$ is the index of the last tangential iteration before $k_{\nu_{1}}$. As in Lemma 3.7, we may now use the assumption that $\Gamma_{k_{\tau_{1}}}\geq\varsigma$ and the monotonicity of the logarithm to obtain that
    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|\leq\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right)\leq\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\sqrt{\frac{2}{\varsigma}}\kappa_{T}\right)=\kappa_{N}.$ (3.23)
    Using the switching condition (2.12), we obtain that, for $\nu\in\{\nu_{0},\ldots,\nu_{1}\}$,
    $\|c_{k_{\nu}}\|\geq\beta\,\alpha_{T,k_{\nu}}\|g_{T,k_{\nu}}\|.$ (3.24)
    Observe now that, using the assumption that $\Gamma_{k_{\nu}}\geq\varsigma$, and (3.17),
    $\alpha_{T,k_{\nu}}=\frac{\eta}{\sqrt{\Gamma_{k_{\nu}}+\varsigma}}\geq\frac{\eta}{\sqrt{2\Gamma_{k_{\nu}}}}\geq\frac{\eta}{\sqrt{2\kappa_{T}}}.$
    Thus, (3.24) yields that, for $\nu\in\{\nu_{0},\ldots,\nu_{1}\}$
    $\|c_{k_{\nu}}\|\ \geq\frac{\beta\eta}{\sqrt{2\kappa_{T}}}\|g_{T,k_{\nu}}\|.$ (3.25)
    With (3.23), this implies that
    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\|g_{T,k_{\nu}}\|\leq\frac{\sqrt{2\kappa_{T}}}{\beta\eta}\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|\leq\frac{\kappa_{N}\,\sqrt{2\kappa_{T}}}{\beta\eta}.$
    Summing this bound with (3.23) gives (3.20). $\Box$
3.7 Combined complexity
The combined complexity may now be derived by assembling the above results.
Theorem 3.9
Suppose that AS.0-AS.7 hold. Then $\frac{1}{k+1}\sum_{j=0}^{k}\Big(\|g_{T,j}\|+\|c_{j}\|\Big)\leq\frac{\kappa_{\mbox{\tiny ADSW,1}}}{\sqrt{k+1}}+\frac{\kappa_{\mbox{\tiny ADSW,2}}}{k+1}={\pazocal{O}}\left(\frac{1}{\sqrt{k+1}}\right),$ (3.26) where $\kappa_{\mbox{\tiny ADSW,1}}=\kappa_{T}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right)+\frac{4\kappa_{N}\,\kappa_{g}}{\beta\eta}\;\;\mbox{ and }\;\;\kappa_{\mbox{\tiny ADSW,2}}=\kappa_{N}\,\left(1+\frac{2\sqrt{2}}{\beta\eta}\right).$
  • Proof.    Now consider iterations from 0 to $k$ of both types (tangential and normal) by setting $\min[k_{\nu_{0}},k_{\tau_{0}}]=0$ and $\max[k_{\nu_{1}},k_{\tau_{1}}]=k$ (as in Lemma 3.5). We then obtain, by combining (3.18) and (3.20), that
    $\sum_{j=0}^{k}$ $\big(\|g_{T,j}\|+\|c_{j}\|\Big)=\sum_{\tau=\tau_{0}}^{\tau_{1}}\Big(\|g_{T,k_{\tau}}\|+\|c_{k_{\tau}}\|\Big)+\sum_{\nu=\nu_{0}}^{\nu_{1}}\Big(\|g_{T,k_{\nu}}\|+\|c_{k_{\nu}}\|\Big)$
    $\leq\kappa_{T}\sqrt{k+1}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right)+\kappa_{N}\,\left(1+\frac{2\sqrt{2}}{\beta\eta}\right),$
    where we used the inequalities $\tau_{1}\leq k_{\tau_{1}}\leq k$ and $k_{\nu_{1}}\leq k$. The bound (3.26) is finally obtained by dividing both sides by $k+1$. $\Box$
4 Numerical illustration
To illustrate the behaviour of ADSWITCH, we coded the algorithm in Matlab and applied it on a set of small dimensional problems from the CUTEst collection [15] as supplied in Matlab by S2MPJ [21]. Our implementation uses the Newton normal step (2.19) and the constants
$\beta=0.01,\;\;\;\;\eta=1,\;\;\;\;\theta=1000,\;\;\;\;\delta=\varsigma=10^{-5}\;\;\mbox{and}\;\;\omega=1.$
For a given $\epsilon\in(0,1)$ a run on a given test problem is deemed successful at iteration $k$ if
$\max[\|g_{T,k}\|,\|c_{k}\|]\leq\epsilon\;\;\mbox{(convergence)}\;\;$ (4.1)
or
$\|J_{k}^{T}c_{k}\|\leq\epsilon\;\;\mbox{but}\;\;\|c_{k}\|>\epsilon\;\;\mbox{(infeasible critical point)}\;\;$ (4.2)
or if a sufficiently optimal function value was found in the sense that
[…]
where $f_{*}$ is the best known feasible function value for the problem. Optimization was also stopped after a maximum of 100000 iterations.
We first ran all test problems with accurate gradient values and $\epsilon=10^{-5}$. The complete results are reported in appendix, and may be broadly summarized as follows.
  1. 1.
    The algorithm appears to converge as predicted. Despite the fact that our current analysis does not cover the case where the Jacobian may lose rank, the algorithm did solve a number of instances where this occurs, sometimes finding an infeasible critical point of the constraints’ violation [^5] This occurs, for instance, at the starting point of HS61..
  2. 2.
    Its performance and reliability is (as could be anticipated) dominated by that of AdaGrad for the tangential step. Since this method is purely first-order, it often performs well, but may fail to solve ill-conditioned problems in a reasonable number of iterations. This is also the case for ADSWITCH. Over the 71 problems in our test set, it solves 44 (62%) of them within 750 iterations, 58 (81%) of them within 100000 iterations and fails on 13. While reliability is not a strong point of AdaGrad and ADSWITCH in the deterministic case, the situation is very different when noise is present, as we show below.
We illustrate a few case of satisfactory convergence in Figures 1 and 2. In these figures, one can clearly see the difference in speed of convergence between tangential steps (AdaGrad-like) and normal steps (Newton), and thus that the overall performance is dominated by that of the first-order method defining the tangential step. This is even clearer when convergence is too slow, as shown in Figure 4 where constraint violation remains very small [^6] HS50 has linear equality constraints only., while very slow convergence of the projected gradient results in a large number of iterations. The left panel of Figure 1 (ORTHREGA with $n=133$ and $m=64$) is also interesting because it shows the switching condition (2.12) in action: the algorithm first approaches an infeasible critical point of the constraints’ violation until the switching rule allows the tangential step to take over, causing the iterates to escape.
Figure 1: Evolution of the optimality measure (left: BYRDSPHR, right: ORTHREGA)
Figure 2: Evolution of the optimality measure (left: LUKVLE2, right: BT1)
Figure 3: Evolution of the optimality measure (HS50)
relative number of number of
noise level total failures total successes
5% 6 61
15% 7 53
25% 7 53
50% 12 51
Figure 4: Reliability of the ADSWITCH algorithm in the presence of relative Gaussian noise on the gradient.
In order to illustrate the resilience of ADSWITCH when noise is present, we also ran the algorithm on all instances of our test problems while adding relative Gaussian noise (of zero mean and unit variance) to the gradient. Each test problem was run ten times independently, and for four levels (5%, 15%, 25% and 50%) of noise. The termination $\epsilon$ was set to $10^{-3}$ in (4.1)–(4.3). The table in Figure 4 reports, for each of these levels, the number of problems for which all ten runs failed (middle column) and the number of problems for which all ten runs converged (rightmost column). This table shows that the reliability of the algorithm is remarkably insensitive to the noise level. It is indeed noteworthy that around two thirds of the considered test problems could be solved with top reliability even if their gradients are perturbed by 50% relative noise, which means that barely one significant digit is correct. This stable behaviour is also visible in Figure 5 showing the performance of the algorithm on the BT1 problem for increasing noise levels.
Figure 5: Evolution of the two components of the optimality measure for a run of the noisy BT1 problem, with 5% (top left), 15% (top right), 25% (bottom left) and 50% (bottom right) relative Gaussian noise on the gradient
5 Conclusions and perspectives
false We have proposed a very simple first-order algorithm for solving nonlinear optimization problems with nonlinear equality constraints. This algorithm adaptively selects steps in the plane tangent to the constraints or steps that reduce infeasibility. It does so without using a merit function or a filter to enforce convergence, but instead relies on the simple switching condition (2.12). The tangent steps are based on the AdaGrad method for unconstrained minimization. As is the case for AdaGrad, the objective function is never evaluated. We have analyzed its worst-case evaluation complexity and obtain global convergence rates which match the best known rates for unconstrained problems. Numerical experiments have been presented indicating that the performance of the algorithm is comparable to that of first-order methods for unconstrained problems.
At this stage, several theoretical questions remain open for further research. These include a stochastic convergence analysis, covering the cases of rank-deficient Jacobians and/or unbounded gradients, the use of alternative minimization methods in the tangential step (such as ADAM [23], ASTR1 [19], OFFAR [20] or more standard unconstrained minimizing techniques using objective function’s values). On the practical side, further numerical experiments are clearly necessary for assessing the true potential of the new method. The handling of inequality constraints is also of obvious interest.
Acknowledgement
Philippe Toint is grateful for the continued and friendly support of the Algo team at Toulouse IRIT. The authors also thank Benedetta Morini and Stefania Bellavia for pointing out a number of typos.
Appendix; Details of the Numerical Results
The following tables give the details of the numerical results discussed in Section 4.
  • Table A.1 reports the results of the noiseless runs. Columns 2 and 3 indicate the number of variables and the number of constraints. The column “#its” gives the number of iterations and the column “exitc” the termination condition (“convg” = convergence, “infeas” = infeasible critical point, “maxit” = maximum number of iterations reached).
  • Tables A.2, A.3, A.4 and A.5 report the average statistics of 10 independent runs for each problem, for increasing relative Gaussian noise levels. The last column of the table gives the number of successful runs (among 10).
Problem $n$ $m$ $f(x)$ $\|g_{T}(x)\|$ $\|c(x)\|$ #its exitc
BT1 2 1 -9.999918e-01 9.64e-07 8.25e-08 141 convg
BT2 3 1 +3.256821e-02 9.34e-07 6.29e-07 186 convg
BT3 5 3 +4.093023e+00 8.58e-07 1.76e-08 70 convg
BT4 3 2 -4.551055e+01 4.49e-07 1.16e-07 16 convg
BT5 3 2 +9.617152e+02 7.00e-07 1.86e-07 46 convg
BT6 5 2 +2.770448e-01 8.75e-07 1.55e-08 107 convg
BT7 5 3 10572 infeas
BT8 5 2 21 infeas
BT9 4 2 -1.000000e+00 9.95e-07 4.92e-07 2481 convg
BT10 2 2 -1.000000e+00 0.00e+00 5.85e-09 6 convg
BT11 5 3 +8.248909e-01 9.55e-07 7.76e-07 351 convg
BT12 5 3 +6.188119e+00 9.46e-07 4.53e-07 240 convg
BYRDSPHR 3 2 -4.683300e+00 8.94e-07 3.33e-08 50 convg
DIXCHLNG 10 5 +2.471898e+03 9.98e-07 2.67e-07 2722 convg
EIGENA2 110 55 +2.448710e-13 9.90e-07 2.06e-09 696 convg
EIGENACO 110 55 +2.436158e-13 9.87e-07 9.41e-16 389 convg
EIGENB2 110 55 +1.800000e+01 6.82e-07 1.78e-15 18 convg
EIGENBCO 110 55 +9.000000e+00 6.82e-07 1.66e-15 18 convg
ELEC 75 25 +2.438128e+02 1.00e-06 4.79e-07 65380 convg
GENHS28 10 8 +9.271737e-01 9.27e-07 1.36e-11 85 convg
HS100LNP 7 2 +6.806301e+02 8.57e-07 2.47e-07 113 convg
HS6 2 1 +9.238528e-13 8.60e-07 8.53e-07 55 convg
HS7 2 1 -1.732051e+00 9.26e-07 1.96e-07 284 convg
HS8 2 2 -1.000000e+00 0.00e+00 1.32e-08 4 convg
HS9 2 1 -5.000000e-01 8.87e-07 7.11e-15 22 convg
HS26 3 1 +1.303916e-09 1.00e-06 5.10e-07 18098 convg
HS27 3 1 +3.999998e-02 9.79e-07 5.23e-07 271 convg
HS28 3 1 +9.687529e-13 9.02e-07 1.33e-15 137 convg
HS39 4 2 -1.000000e+00 9.95e-07 4.92e-07 2481 convg
HS40 4 3 -2.500002e-01 9.74e-07 3.24e-07 600 convg
HS42 4 2 +1.385786e+01 9.85e-07 8.75e-07 92 convg
HS46 5 2 +8.808661e-09 1.00e-06 2.35e-07 13907 convg
HS47 5 3 +4.650485e-10 1.00e-06 6.88e-07 26784 convg
HS48 5 2 +2.912966e-13 9.31e-07 2.81e-15 177 convg
HS50 5 3 +2.888537e-13 9.99e-07 6.63e-14 11166 convg
HS51 5 3 +8.485715e-14 5.67e-07 1.46e-15 19 convg
HS52 5 3 +5.326648e+00 9.87e-07 1.76e-09 150 convg
HS61 3 2 2 infeas
HS77 5 2 +2.415051e-01 9.27e-07 1.18e-07 103 convg
HS78 5 3 -2.919700e+00 9.31e-07 1.97e-08 95 convg
HS79 5 3 +7.877683e-02 9.71e-07 2.29e-07 357 convg
LUKVLE1 20 18 +6.232459e+00 9.11e-07 1.20e-08 98 convg
LUKVLE2 20 13 +4.372199e+02 9.93e-07 3.19e-08 604 convg
LUKVLE3 20 2 +2.758657e+01 9.94e-07 8.67e-07 2319 convg
LUKVLE4 20 9 +1.060660e+02 1.36e+02 3.15e-07 100000 maxit
LUKVLE6 21 10 +1.120010e+03 2.10e-01 7.64e-07 100000 maxit
LUKVLE7 20 4 -5.352208e+00 8.90e-07 3.91e-08 111 convg
LUKVLE8 20 18 +2.099621e+03 8.65e-07 3.18e-09 55 convg
LUKVLE9 20 6 +6.560168e+00 1.51e+01 8.57e-07 100000 maxit
LUKVLE10 20 18 +6.639332e+00 7.23e-07 3.27e-07 40 convg
LUKVLE11 18 10 +5.359296e-07 3.53e-05 1.69e-09 100000 maxit
LUKVLE12 17 12 +2.286872e+02 4.44e+01 2.23e+00 100000 maxit
LUKVLE13 18 10 +5.483076e+01 4.24e-06 1.20e-08 100000 maxit
LUKVLE14 18 10 +4.180657e+04 7.86e-01 7.02e-07 100000 maxit
LUKVLE15 17 12 +7.527144e+01 1.14e+01 9.10e-07 100000 maxit
LUKVLE16 17 12 +8.040898e+01 2.71e+01 3.79e+00 100000 maxit
LUKVLE17 17 12 +1.190180e+02 5.48e+01 8.37e+00 100000 maxit
LUKVLE18 17 12 +2.627594e+01 1.71e+01 8.37e+00 100000 maxit
LUKVLI4 20 9 +1.060660e+02 1.36e+02 3.15e-07 100000 maxit
MARATOS 2 1 -1.000000e+00 9.27e-07 2.26e-07 168 convg
MWRIGHT 5 3 +2.497881e+01 9.23e-07 1.04e-08 114 convg
ORTHRDM2 9 3 +1.087113e-13 6.51e-07 7.14e-07 31 convg
ORTHRDS2 9 3 +5.921433e-14 4.81e-07 2.63e-08 33 convg
ORTHREGA 133 64 +3.503002e+02 9.82e-07 1.66e-08 705 convg
ORTHREGB 27 6 +3.057742e-14 3.50e-07 1.02e-08 15 convg
ORTHREGC 15 5 +1.125393e-12 9.24e-07 7.22e-08 50 convg
ORTHREGD 43 20 +2.179046e+02 1.74e+01 1.09e-03 100000 maxit
ORTHRGDM 43 20 +5.067350e+01 8.20e+00 1.97e-03 100000 maxit
ORTHRGDS 43 20 +6.210193e+00 9.97e-07 8.69e-07 474 convg
S316m322 2 1 0 infeas
SPINOP 11 9 +8.435060e-02 7.91e-04 2.20e-06 100000 maxit
Table A.1: Results of running ADSWITCH in the abscence of noise
Problem $n$ $m$ avr $f(x)$ avr $\|g_{T}(x)\|$ avr $\|c(x)\|$ avr #its #success
BT1 3 1 -5.681296e-01 7.73e-04 3.20e-04 8.92e+01 10
BT2 5 3 +3.257271e-02 1.51e-03 3.49e-04 1.18e+02 10
BT3 3 2 +4.089376e+00 5.05e-02 6.34e-04 2.50e+04 10
BT4 3 2 -4.551215e+01 1.87e-02 5.03e-04 8.88e+01 10
BT5 5 2 +9.617275e+02 1.60e-01 9.36e-05 1.10e+03 10
BT6 5 3 +2.770621e-01 9.19e-03 2.69e-04 3.50e+03 10
BT7 5 2 2.51e+03 10
BT8 4 2 +1.000238e+00 6.34e-04 3.41e-04 1.47e+01 10
BT9 2 2 -1.000535e+00 9.87e-04 5.37e-04 1.15e+03 10
BT10 5 3 -1.000038e+00 0.00e+00 5.91e-05 5.00e+00 10
BT11 5 3 +8.246795e-01 4.64e-02 4.07e-04 1.29e+04 10
BT12 3 2 +6.188179e+00 6.06e-03 6.49e-04 1.29e+02 10
BYRDSPHR 10 5 -4.683084e+00 1.21e-02 1.63e-05 1.85e+02 10
DIXCHLNG 110 55 +2.471678e+03 1.25e+01 3.47e-04 9
EIGENA2 110 55 +2.507664e-07 1.00e-03 4.03e-04 3.98e+02 10
EIGENACO 110 55 +2.410396e-07 9.81e-04 9.35e-16 2.31e+02 10
EIGENB2 110 55 +3.530739e-06 1.00e-03 6.16e-04 1.10e+03 10
EIGENBCO 75 25 +5.356264e-06 9.58e-04 5.29e-04 1.07e+03 10
ELEC 10 8 +2.438048e+02 1.86e-01 4.33e-04 1.00e+05 10
GENHS28 7 2 +9.274072e-01 2.82e-02 6.75e-06 3.01e+03 10
HS100LNP 2 1 +6.806369e+02 5.73e-01 1.91e-04 1
HS6 2 1 +9.637182e-07 8.72e-04 2.76e-04 2.96e+01 10
HS7 2 2 -1.732173e+00 9.72e-04 4.26e-04 1.65e+02 10
HS8 2 1 -1.000000e+00 0.00e+00 1.32e-08 4.00e+00 10
HS9 3 1 -4.998315e-01 3.11e-03 5.86e-15 1.41e+01 10
HS26 3 1 +1.374024e-05 1.01e-03 7.19e-04 2.19e+02 10
HS27 3 1 +3.999576e-02 9.84e-04 2.58e-04 6.39e+01 10
HS28 4 2 +1.022606e-06 9.25e-04 6.55e-16 7.09e+01 10
HS39 4 3 -1.000759e+00 9.86e-04 7.61e-04 1.08e+03 10
HS40 4 2 -2.501555e-01 1.42e-02 4.17e-04 2.71e+02 10
HS42 5 2 +1.386050e+01 1.79e-01 2.38e-04 9.08e+02 10
HS46 5 3 +4.853021e-06 9.96e-04 5.22e-04 1.31e+02 10
HS47 5 2 +1.474970e-05 1.03e-03 4.84e-04 7.36e+02 10
HS48 5 3 +3.056976e-07 9.53e-04 1.88e-15 9.29e+01 10
HS50 5 3 +3.332153e-07 1.07e-03 1.07e-13 6.29e+03 10
HS51 5 3 +1.176429e-07 6.55e-04 1.69e-15 1.03e+01 10
HS52 3 2 +5.326494e+00 1.14e-01 6.34e-05 3.00e+04 10
HS61 5 2 1.00e+00 10
HS77 5 3 +2.415176e-01 1.14e-02 3.55e-04 2.04e+03 10
HS78 5 3 -2.918759e+00 6.58e-02 3.58e-04 6
HS79 20 18 +7.879160e-02 2.26e-03 3.19e-04 1.24e+02 10
LUKVLE1 20 13 +6.234044e+00 2.71e-03 3.84e-04 1.13e+02 10
LUKVLE2 20 2 +4.371168e+02 3.46e+00 5.67e-04 1.00e+05 10
LUKVLE3 20 9 +2.757471e+01 1.64e-01 6.39e-04 1.00e+05 10
LUKVLE4 21 10 +1.058513e+02 1.35e+02 5.50e-06 1.00e+05 10
LUKVLE6 20 4 +1.119942e+03 4.69e-01 5.08e-04 1.00e+05 10
LUKVLE7 20 18 -5.353248e+00 1.18e-01 3.97e-04 0
LUKVLE8 20 6 1.13e+04 10
LUKVLE9 20 18 +6.555645e+00 1.51e+01 1.66e-04 1.00e+05 10
LUKVLE10 18 10 +6.639743e+00 1.44e-02 2.76e-04 4.84e+03 10
LUKVLE11 17 12 +3.438763e-05 1.09e-03 4.80e-04 6.96e+03 10
LUKVLE12 18 10 +2.286872e+02 4.44e+01 2.23e+00 1.00e+05 10
LUKVLE13 18 10 +5.483466e+01 8.14e-02 4.72e-04 1.00e+05 10
LUKVLE14 17 12 +4.180658e+04 8.46e-01 7.29e-04 1.00e+05 10
LUKVLE15 17 12 +7.511104e+01 1.14e+01 1.81e-04 0
LUKVLE16 17 12 +8.040898e+01 2.71e+01 3.79e+00 0
LUKVLE17 17 12 +1.190180e+02 5.48e+01 8.37e+00 0
LUKVLE18 20 9 +2.627594e+01 1.71e+01 8.37e+00 0
LUKVLI4 2 1 +1.063358e+02 1.38e+02 5.60e-06 1.00e+05 10
MARATOS 5 3 -1.000295e+00 9.53e-04 5.91e-04 9.98e+01 10
MWRIGHT 9 3 +2.497543e+01 1.45e-01 4.88e-04 9
ORTHRDM2 9 3 +1.034714e-07 6.19e-04 3.51e-04 2.16e+01 10
ORTHRDS2 133 64 +8.522014e-08 5.48e-04 3.05e-04 2.14e+01 10
ORTHREGA 27 6 +3.502947e+02 6.49e-02 7.05e-04 1.00e+05 10
ORTHREGB 15 5 +5.311601e-08 3.99e-04 3.80e-05 1.06e+01 10
ORTHREGC 43 20 +6.775424e-07 8.16e-04 5.67e-04 2.69e+01 10
ORTHREGD 43 20 2.13e+02 10
ORTHRGDM 43 20 1.86e+03 10
ORTHRGDS 2 1 +6.210324e+00 2.21e-02 5.60e-04 0
S316m322 11 9 0.00e+00 10
SPINOP 11 9 +9.941025e-02 1.10e-03 5.74e-04 6.14e+04 10
Table A.2: Results of running ADSWITCH with 5% relative Gaussian noise
Problem $n$ $m$ avr $f(x)$ avr $\|g_{T}(x)\|$ avr $\|c(x)\|$ avr #its #success
BT1 3 1 -1.720654e-01 2.87e-04 2.80e-04 2.19e+02 10
BT2 5 3 +3.257574e-02 2.86e-03 5.14e-04 1.45e+02 10
BT3 3 2 +4.089932e+00 6.74e-02 6.34e-04 7.06e+04 10
BT4 3 2 -4.551083e+01 5.03e-02 1.67e-04 3.36e+02 10
BT5 5 2 +9.617453e+02 2.74e-01 4.39e-05 4.57e+03 10
BT6 5 3 +2.771418e-01 1.71e-02 3.62e-04 7
BT7 5 2 1.97e+03 10
BT8 4 2 +1.000238e+00 3.77e-04 3.41e-04 1.38e+01 10
BT9 2 2 -1.000468e+00 7.93e-04 4.70e-04 6.71e+02 10
BT10 5 3 -1.000038e+00 0.00e+00 5.91e-05 5.00e+00 10
BT11 5 3 +8.248886e-01 6.47e-02 4.69e-04 7
BT12 3 2 +6.188227e+00 1.06e-02 4.39e-04 1.90e+02 10
BYRDSPHR 10 5 -4.680516e+00 3.82e-02 1.76e-04 4.58e+02 10
DIXCHLNG 110 55 +2.471747e+03 3.29e+01 5.13e-04 8
EIGENA2 110 55 +2.864156e-07 1.07e-03 5.67e-04 3.97e+02 10
EIGENACO 110 55 +2.500993e-07 9.99e-04 9.37e-16 2.35e+02 10
EIGENB2 110 55 +2.989956e-06 1.02e-03 2.20e-04 1.05e+03 10
EIGENBCO 75 25 +1.674676e-06 8.43e-04 4.82e-04 1.35e+03 10
ELEC 10 8 +2.438072e+02 4.00e-01 5.36e-04 6
GENHS28 7 2 +9.275303e-01 3.38e-02 6.75e-06 3.59e+04 10
HS100LNP 2 1 +6.806511e+02 1.04e+00 3.81e-04 0
HS6 2 1 +8.822774e-07 8.34e-04 5.77e-04 2.94e+01 10
HS7 2 2 -1.732173e+00 9.28e-04 4.25e-04 1.45e+02 10
HS8 2 1 -1.000000e+00 0.00e+00 1.32e-08 4.00e+00 10
HS9 3 1 -4.993365e-01 7.17e-03 6.76e-15 2.20e+01 10
HS26 3 1 +1.632328e-05 1.15e-03 4.97e-04 2.15e+02 10
HS27 3 1 +3.999706e-02 9.14e-04 2.16e-04 4.66e+01 10
HS28 4 2 +1.738691e-06 1.20e-03 7.54e-16 6.87e+01 10
HS39 4 3 -1.000411e+00 1.07e-03 4.13e-04 5.98e+02 10
HS40 4 2 -2.493358e-01 3.90e-02 4.84e-06 4.18e+02 10
HS42 5 2 +1.386567e+01 2.88e-01 1.42e-04 1.38e+03 10
HS46 5 3 +7.154893e-06 1.06e-03 3.65e-04 1.24e+02 10
HS47 5 2 +1.763663e-05 1.17e-03 5.01e-04 6.97e+02 10
HS48 5 3 +3.810065e-07 1.05e-03 1.86e-15 9.47e+01 10
HS50 5 3 +4.796120e-07 1.29e-03 9.03e-14 6.20e+03 10
HS51 5 3 +1.153904e-07 6.35e-04 1.57e-15 9.90e+00 10
HS52 3 2 +5.326745e+00 1.30e-01 6.34e-05 5
HS61 5 2 1.00e+00 10
HS77 5 3 +2.415599e-01 1.79e-02 2.39e-04 5.11e+04 10
HS78 5 3 -2.918552e+00 6.92e-02 4.09e-04 4
HS79 20 18 +7.879634e-02 2.28e-03 5.15e-04 1.54e+02 10
LUKVLE1 20 13 +6.233718e+00 7.71e-03 3.12e-04 1.20e+02 10
LUKVLE2 20 2 +4.371783e+02 7.13e+00 4.04e-04 8
LUKVLE3 20 9 +2.757958e+01 5.72e-01 4.40e-04 9
LUKVLE4 21 10 +1.063795e+02 1.41e+02 8.44e-06 1.00e+05 10
LUKVLE6 20 4 +1.119853e+03 1.16e+00 5.46e-04 1.00e+05 10
LUKVLE7 20 18 -5.350163e+00 4.22e-01 2.45e-04 0
LUKVLE8 20 6 1.12e+03 10
LUKVLE9 20 18 +6.707674e+00 1.60e+01 2.84e-04 1.00e+05 10
LUKVLE10 18 10 +6.473732e+00 2.91e-02 3.32e-04 1.19e+04 10
LUKVLE11 17 12 +2.849669e-01 5.36e-03 5.57e-04 9
LUKVLE12 18 10 +2.286872e+02 4.44e+01 2.23e+00 1.00e+05 10
LUKVLE13 18 10 +5.483592e+01 1.08e-01 5.89e-04 1.00e+05 10
LUKVLE14 17 12 +4.180658e+04 1.02e+00 3.89e-04 1.00e+05 10
LUKVLE15 17 12 +7.514427e+01 1.15e+01 1.81e-04 0
LUKVLE16 17 12 +8.040898e+01 2.71e+01 3.79e+00 0
LUKVLE17 17 12 +1.190180e+02 5.48e+01 8.37e+00 0
LUKVLE18 20 9 +2.627594e+01 1.71e+01 8.37e+00 0
LUKVLI4 2 1 +1.094837e+02 1.56e+02 6.13e-06 1.00e+05 10
MARATOS 5 3 -1.000232e+00 7.86e-04 4.65e-04 9.35e+01 10
MWRIGHT 9 3 +2.497649e+01 2.93e-01 5.48e-04 7
ORTHRDM2 9 3 +9.737407e-08 5.56e-04 4.05e-04 2.27e+01 10
ORTHRDS2 133 64 +1.610065e-07 7.37e-04 4.02e-04 2.24e+01 10
ORTHREGA 27 6 +3.502994e+02 1.19e-01 4.88e-04 5
ORTHREGB 15 5 +7.691602e-08 4.91e-04 8.40e-05 1.09e+01 10
ORTHREGC 43 20 +9.146250e-07 8.78e-04 3.23e-04 2.67e+01 10
ORTHREGD 43 20 1.41e+02 10
ORTHRGDM 43 20 1.66e+03 10
ORTHRGDS 2 1 +6.210852e+00 5.11e-02 5.33e-04 0
S316m322 11 9 0.00e+00 10
SPINOP 11 9 +1.134528e-01 1.43e-03 6.21e-04 4.12e+04 10
Table A.3: Results of running ADSWITCH with 15% relative Gaussian noise
Problem $n$ $m$ avr $f(x)$ avr $\|g_{T}(x)\|$ avr $\|c(x)\|$ avr #its #success
BT1 3 1 -1.628888e-01 1.65e-05 3.72e-04 1.86e+02 10
BT2 5 3 +3.257851e-02 5.28e-03 3.90e-04 2.48e+02 10
BT3 3 2 +4.089948e+00 6.60e-02 6.34e-04 9.29e+04 10
BT4 3 2 -4.551030e+01 1.04e-01 2.53e-04 5.51e+02 10
BT5 5 2 +9.617367e+02 2.09e-01 1.93e-04 6.10e+03 10
BT6 5 3 +2.771881e-01 2.53e-02 6.25e-04 1
BT7 5 2 1.51e+03 10
BT8 4 2 +1.000238e+00 2.43e-04 3.41e-04 1.42e+01 10
BT9 2 2 -1.000286e+00 6.60e-04 2.91e-04 4.16e+02 10
BT10 5 3 -1.000038e+00 0.00e+00 5.91e-05 5.00e+00 10
BT11 5 3 +8.259372e-01 1.09e-01 4.43e-04 6
BT12 3 2 +6.188639e+00 2.01e-02 3.28e-04 2.05e+02 10
BYRDSPHR 10 5 -4.672953e+00 8.00e-02 9.61e-05 5.88e+02 10
DIXCHLNG 110 55 +2.471810e+03 3.74e+01 4.88e-04 6
EIGENA2 110 55 +2.814117e-07 1.06e-03 3.87e-04 4.03e+02 10
EIGENACO 110 55 +2.651002e-07 1.03e-03 9.35e-16 2.43e+02 10
EIGENB2 110 55 +2.653621e-06 1.02e-03 5.90e-04 1.27e+03 10
EIGENBCO 75 25 +6.332410e-07 8.51e-04 4.14e-04 1.56e+03 10
ELEC 10 8 +2.438104e+02 5.26e-01 5.69e-04 6
GENHS28 7 2 +9.281926e-01 5.42e-02 6.75e-06 2.64e+04 10
HS100LNP 2 1 +6.806694e+02 1.50e+00 6.92e-04 0
HS6 2 1 +9.704023e-07 8.32e-04 4.74e-04 3.12e+01 10
HS7 2 2 -1.732196e+00 7.94e-04 5.04e-04 1.51e+02 10
HS8 2 1 -1.000000e+00 0.00e+00 1.32e-08 4.00e+00 10
HS9 3 1 -4.981924e-01 1.09e-02 2.06e-14 7.99e+01 10
HS26 3 1 +1.680639e-05 1.18e-03 5.93e-04 2.00e+02 10
HS27 3 1 +4.000247e-02 9.90e-04 1.26e-04 4.02e+01 10
HS28 4 2 +2.764455e-06 1.42e-03 8.77e-16 7.06e+01 10
HS39 4 3 -1.000434e+00 7.07e-04 4.45e-04 3.96e+02 10
HS40 4 2 -2.232808e-01 6.51e-02 1.23e-04 5.08e+02 10
HS42 5 2 +1.387811e+01 4.89e-01 3.55e-04 3.79e+03 10
HS46 5 3 +7.617645e-06 1.14e-03 6.62e-04 1.16e+02 10
HS47 5 2 +2.249967e-05 1.37e-03 3.85e-04 6.16e+02 10
HS48 5 3 +5.445390e-07 1.25e-03 1.39e-15 8.88e+01 10
HS50 5 3 +7.538284e-07 1.61e-03 9.79e-14 6.08e+03 10
HS51 5 3 +1.182718e-07 6.25e-04 1.67e-15 1.05e+01 10
HS52 3 2 +5.327150e+00 1.37e-01 6.34e-05 1
HS61 5 2 1.00e+00 10
HS77 5 3 +2.416327e-01 2.47e-02 2.62e-04 8
HS78 5 3 -2.918216e+00 8.49e-02 3.36e-04 1
HS79 20 18 +7.879582e-02 3.00e-03 6.26e-04 2.09e+02 10
LUKVLE1 20 13 +4.363520e+00 6.83e-03 3.41e-04 1.25e+02 10
LUKVLE2 20 2 +4.371531e+02 5.88e+00 6.31e-04 9
LUKVLE3 20 9 +2.758052e+01 1.14e+00 6.23e-04 7
LUKVLE4 21 10 +1.146093e+02 1.73e+02 6.14e-06 1.00e+05 10
LUKVLE6 20 4 +1.119798e+03 2.06e+00 4.77e-04 1.00e+05 10
LUKVLE7 20 18 -5.349869e+00 4.53e-01 3.32e-04 0
LUKVLE8 20 6 3.95e+02 10
LUKVLE9 20 18 +6.467339e+00 1.45e+01 1.53e-04 1.00e+05 10
LUKVLE10 18 10 +6.464724e+00 2.88e-02 5.32e-04 4.79e+04 10
LUKVLE11 17 12 +6.422930e-05 1.84e-03 3.02e-04 5.17e+03 10
LUKVLE12 18 10 +2.286872e+02 4.44e+01 2.23e+00 1.00e+05 10
LUKVLE13 18 10 +5.501071e+01 1.89e-01 3.69e-04 1.00e+05 10
LUKVLE14 17 12 +4.180664e+04 1.77e+00 4.53e-04 1.00e+05 10
LUKVLE15 17 12 +7.384004e+01 1.14e+01 1.82e-04 0
LUKVLE16 17 12 +8.040898e+01 2.71e+01 3.79e+00 0
LUKVLE17 17 12 +1.190180e+02 5.48e+01 8.37e+00 0
LUKVLE18 20 9 +2.627594e+01 1.71e+01 8.37e+00 0
LUKVLI4 2 1 +1.094630e+02 1.55e+02 9.07e-06 1.00e+05 10
MARATOS 5 3 -9.998270e-01 8.83e-03 2.92e-04 3.76e+01 10
MWRIGHT 9 3 +2.497751e+01 2.90e-01 5.26e-04 7
ORTHRDM2 9 3 +1.408678e-07 6.56e-04 4.15e-04 2.25e+01 10
ORTHRDS2 133 64 +1.867325e-07 7.67e-04 4.32e-04 2.25e+01 10
ORTHREGA 27 6 +3.503036e+02 1.71e-01 4.19e-04 3
ORTHREGB 15 5 +6.680941e-08 4.66e-04 2.55e-04 1.21e+01 10
ORTHREGC 43 20 +5.552793e-07 7.69e-04 3.77e-04 2.81e+01 10
ORTHREGD 43 20 1.44e+02 10
ORTHRGDM 43 20 1.60e+03 10
ORTHRGDS 2 1 +6.211177e+00 6.17e-02 6.38e-04 0
S316m322 11 9 0.00e+00 10
SPINOP 11 9 +1.331340e-01 1.97e-03 5.60e-04 2.56e+04 10
Table A.4: Results of running ADSWITCH with 25% relative Gaussian noise
Problem $n$ $m$ avr $f(x)$ avr $\|g_{T}(x)\|$ avr $\|c(x)\|$ avr #its #success
BT1 3 1 -8.595453e-01 9.99e-02 4.59e-04 2.67e+02 10
BT2 5 3 +3.258385e-02 7.70e-03 4.76e-04 7.86e+02 10
BT3 3 2 +4.090844e+00 1.06e-01 6.34e-04 9
BT4 3 2 -4.550471e+01 2.77e-01 2.28e-04 1.19e+03 10
BT5 5 2 +9.618963e+02 6.57e-01 2.61e-04 5.57e+03 10
BT6 5 3 +2.773363e-01 2.99e-02 5.98e-04 0
BT7 5 2 1.14e+03 10
BT8 4 2 +1.000243e+00 1.98e-03 3.41e-04 1.46e+01 10
BT9 2 2 -1.000035e+00 1.17e-02 2.21e-04 1.98e+02 10
BT10 5 3 -1.000038e+00 0.00e+00 5.91e-05 5.00e+00 10
BT11 5 3 +8.258014e-01 1.25e-01 5.58e-04 5
BT12 3 2 +6.189020e+00 2.82e-02 3.48e-04 5.23e+02 10
BYRDSPHR 10 5 -4.664093e+00 1.11e-01 1.19e-04 8.67e+02 10
DIXCHLNG 110 55 +2.472434e+03 6.46e+01 5.25e-04 5
EIGENA2 110 55 +9.973427e-07 1.86e-03 3.57e-04 4.15e+02 10
EIGENACO 110 55 +2.948147e-07 1.08e-03 9.41e-16 2.81e+02 10
EIGENB2 110 55 +3.675137e-06 1.07e-03 5.37e-04 8.56e+02 10
EIGENBCO 75 25 +2.125972e-06 7.45e-04 5.15e-04 1.94e+03 10
ELEC 10 8 +2.438199e+02 6.84e-01 5.62e-04 2
GENHS28 7 2 +9.280133e-01 5.07e-02 6.75e-06 1
HS100LNP 2 1 +6.806891e+02 1.84e+00 8.94e-04 0
HS6 2 1 +1.921030e-03 1.60e-02 3.73e-04 2.96e+01 10
HS7 2 2 -1.732161e+00 8.73e-04 3.80e-04 9.28e+01 10
HS8 2 1 -1.000000e+00 0.00e+00 1.32e-08 4.00e+00 10
HS9 3 1 -4.968118e-01 1.36e-02 1.10e-14 7.22e+01 10
HS26 3 1 +3.827215e-05 2.15e-03 4.57e-04 1.67e+02 10
HS27 3 1 +4.007659e-02 2.90e-03 2.31e-04 9.16e+01 10
HS28 4 2 +2.028082e-05 2.71e-03 8.44e-16 7.52e+01 10
HS39 4 3 -1.000374e+00 5.82e-03 4.58e-04 1.94e+02 10
HS40 4 2 -1.981579e-01 4.81e-02 1.35e-04 3.81e+02 10
HS42 5 2 +1.391016e+01 5.52e-01 2.13e-04 4.87e+03 10
HS46 5 3 +4.013015e-05 1.67e-03 5.27e-04 1.56e+02 10
HS47 5 2 +5.386687e-05 2.46e-03 5.13e-04 4.49e+02 10
HS48 5 3 +8.988756e-06 4.31e-03 1.39e-15 8.36e+01 10
HS50 5 3 +8.049833e-05 1.25e-02 1.08e-13 5.38e+03 10
HS51 5 3 +2.340881e-07 9.39e-04 1.98e-15 1.68e+01 10
HS52 3 2 +5.328581e+00 1.57e-01 6.34e-05 0
HS61 5 2 1.00e+00 10
HS77 5 3 +2.416904e-01 3.22e-02 4.66e-04 0
HS78 5 3 -2.914874e+00 1.41e-01 2.74e-04 0
HS79 20 18 +7.880086e-02 4.41e-03 5.19e-04 9.58e+02 10
LUKVLE1 20 13 +3.740180e+00 1.06e-02 2.27e-04 1.11e+02 10
LUKVLE2 20 2 +4.371965e+02 8.04e+00 5.47e-04 6
LUKVLE3 20 9 +2.758515e+01 1.13e+00 4.96e-04 5
LUKVLE4 21 10 +1.859096e+02 2.56e+02 9.65e-05 1.00e+05 10
LUKVLE6 20 4 +1.119752e+03 4.22e+00 5.48e-04 1.00e+05 10
LUKVLE7 20 18 -5.347145e+00 5.86e-01 4.82e-04 0
LUKVLE8 20 6 1.90e+02 10
LUKVLE9 20 18 +7.100923e+00 1.90e+01 3.53e-04 1.00e+05 10
LUKVLE10 18 10 +6.459921e+00 2.75e-02 4.97e-04 9.69e+04 10
LUKVLE11 17 12 +1.000256e-04 2.75e-03 5.29e-04 3.72e+03 10
LUKVLE12 18 10 +2.286872e+02 4.44e+01 2.23e+00 1.00e+05 10
LUKVLE13 18 10 +5.501757e+01 3.78e-01 3.38e-04 1.00e+05 10
LUKVLE14 17 12 +4.180692e+04 4.10e+00 3.72e-04 1.00e+05 10
LUKVLE15 17 12 +7.963970e+01 1.27e+01 1.81e-04 0
LUKVLE16 17 12 +8.040898e+01 2.71e+01 3.79e+00 0
LUKVLE17 17 12 +1.190180e+02 5.48e+01 8.37e+00 0
LUKVLE18 20 9 +2.627594e+01 1.71e+01 8.37e+00 0
LUKVLI4 2 1 +1.641033e+02 2.16e+02 2.26e-04 1.00e+05 10
MARATOS 5 3 -1.000135e+00 4.39e-03 3.68e-04 3.51e+01 10
MWRIGHT 9 3 +2.498269e+01 4.85e-01 4.89e-04 4
ORTHRDM2 9 3 +1.938152e-07 7.12e-04 2.85e-04 2.68e+01 10
ORTHRDS2 133 64 +2.075502e-07 7.50e-04 2.86e-04 2.67e+01 10
ORTHREGA 27 6 +3.503053e+02 2.04e-01 6.55e-04 0
ORTHREGB 15 5 +1.903767e-07 7.21e-04 2.66e-04 1.51e+01 10
ORTHREGC 43 20 +7.509051e-07 8.31e-04 5.59e-04 3.67e+01 10
ORTHREGD 43 20 1.65e+02 10
ORTHRGDM 43 20 1.84e+03 10
ORTHRGDS 2 1 +6.212538e+00 9.40e-02 7.00e-04 0
S316m322 11 9 0.00e+00 10
SPINOP 11 9 +3.658931e-01 1.59e-02 4.10e-04 1.53e+03 10
Table A.5: Results of running ADSWITCH with 50% relative Gaussian noise

### v2

@@
 Abstract
-A very simple first-order algorithm is proposed for solving nonlinear optimization problems with nonlinear equality constraints.
+A very simple first-order algorithm is proposed for solving potentially stochastic nonlinear optimization problems with deterministic nonlinear equality constraints.
 This algorithm adaptively selects steps in the plane tangent to the constraints or steps that reduce infeasibility, without using a merit function or filter.
@@
 The objective function is never evaluated by the algorithm, making it suitable for noisy problems.
-Its worst-case evaluation complexity is analyzed, yielding a global convergence rate in ${\pazocal{O}}(1/\sqrt{k})$, which matches the best known rate of first-order methods for unconstrained problems.
+Its worst-case evaluation complexity and convergence is analyzed both in the deterministic and stochastic cases, yielding a global convergence rate in ${\pazocal{O}}(1/\sqrt{k})$ for the deterministic case and in ${\pazocal{O}}(1/k^{1/4})$ for the stochastic one, which match the best known rates of first order methods for unconstrained problems.
 Numerical experiments are presented suggesting that the performance of the algorithm is comparable to that of first-order methods for unconstrained problems, and that its reliability is remarkably stable in the presence of noise on the gradient.
-Keywords: Equality constrained optimization, objective-function-free optimization (OFFO), first-order methods, AdaGrad, evaluation complexity.
+Keywords: Equality constrained optimization, objective-function-free optimization (OFFO), first-order methods, AdaGrad, evaluation complexity, stochastic analysis.
 1 Introduction
 Solving nonlinear problems with nonlinear equality constraints is a central part of more general constrained optimization, and the purpose of this paper is to contribute to the subject by proposing a very simple algorithm for the case where the constraints’ Jacobian is full-rank.
-We introduce a method drawing its inspiration both from the “trust-funnel” approaches [17, 8] and the objective-function free (OFFO) first-order algorithms that have become very popular because of their successful use in deep learning applications.
+We introduce a method drawing its inspiration both from the “trust-funnel” approaches [15, 7] and the objective-function free (OFFO) first-order algorithms that have become very popular because of the successful use in deep learning applications.
 Trust-funnel methods themselves exploit the older idea of decomposing the step computed within an optimization algorithm for constrained problems into a “tangential” and a “normal” components.
-Advocated in particular by Byrd and Omojokun [24] for sequential quadratic programming, this last idea is to consider that a useful minimization step should, on the one hand, improve the value of the objective function without deteriorating the constraints’s violation too much (this is the tangential step, because it can be interpreted as a step (nearly) tangent to the constraint’s surface) and, on the other hand, improve feasibility (the normal step, mostly orthogonal to the constraints’ surface).
+Advocated in particular by Byrd and Omojokun [22] for sequential quadratic programming, the last idea is to consider that a useful minimization step should, one one hand, improve the value of the objective function without deteriorating the constraints’s violation too much (this is the tangential step, because it can be interpreted as a step (nearly) tangent to the constraint’s surface) and, on the other hand, improve feasibility (the normal step, mostly orthogonal to the constraints’ surface).
 This technique has been successfully used by many algorithms, and in particular by trust-funnel ones.
@@
 Such methods are attractive because they avoid computing the value of the objective function entirely (they only rely on gradients).
-This feature makes them quite robust in the presence of noise on the gradient, as is commonly the case when its evaluation involves subsampling.
-Again, our new method exploits this OFFO [^1] Objective-Function-Free Optimization.
+This feature makes them quite robust in the presence of noise on the gradient, as is commonly the case its evluation involves subsampling.
+Again, our new method exploits this OFFO [^1] Objectibe-Function-Free Optimization.
 strategy to achieve robustness.
@@
 Other OFFO methods for constrained optimization have been proposed.
-The stochastic algorithm proposed in [1] uses a combined tangential and normal step for a quadratic model of a merit function.
-A (nontrivial) stochastic analysis of its convergence (albeit not its worst-case complexity) is provided.
-It also allows the use of second-order information, at the cost of a somewhat complicated algorithm.
-[7] extends this proposal to handle inequality constraints and provides an updated convergence analysis.
-The algorithm of [11] is more similar to ours, but differs in significant aspects.
-Although it uses tangential and normal steps, it does so at every iteration.
-The step in the Jacobian’s nullspace is computed by a standard trust-region technique and a special rule is provided to update its radius.
-It also relies on an explicit penalty parameter with is updated in the course of the algorithm.
-Convergence of the algorithm is also analyzed in a stochastic setting.
-The method of [4] is inspired by filter methods (another class of methods for nonlinear optimization [13, 14, 12]) and also alternates between steps which attempts to reduce constraints’ infeasibilities and “optimality” steps improving the objective function value, albeit ignoring the presence of constraints in the second case.
-Good numerical results are reported on deep learning problems, but no convergence theory or complexity analysis is provided [^2] Since the optimality steps ignore the constraints’ geometry, it seems possible that their effect might significantly decrease feasibility and that cycling could occur..
+The stochastic SQP method of [6] is more general in that it also handles inequality constraints and provides a (nontrivial) stochastic analysis of its convergence (albeit not its worst-case complexity).
+It also allow the use of second-order information, at the cost of a quite complicated algorithm.
+The method of [3] is inspired by filter methods (another class of methods for nonlinear optimization [11, 12, 10]) and also alternates between steps which attempts to reduce constraints’s infeasibilities and “optimality” steps improving the objective function value, albeit ignoring the presence of constraints in the second case.
+Good numerical results are reported on deep learning problems, but no convergence theory or complexity analysis is provided [^2] Since the optimality steps ignore the constraints’s geometry, it seems possible that their effect might significantly decrease feasibility and that cycling could occur..
 In view of these comments, we summarize our contributions as follows.
@@
   2. 2.
-    We provide a complete worst-case complexity analysis for the case where the objective’s gradient and the constraints are deterministic, under assumptions similar to those used in [1, 7].
+    We provide a complete worst-case complexity analysis for the case where the objective’s gradient can be deterministic or contaminated by random noise and the constraints are deterministic, under assumptions similar to those used in [6].
   3. 3.
-    We illustrate the practical behaviour of the method on a subset of the CUTEst problems [15] as provided by the S2MPJ environment [21].
+    We illustrate the practical behaviour of the method on a subset of the CUTEst problems [13] as provided by the S2MPJ environment [19], both in the deterministic and noisy contexts.
 The paper is organized as follows.
@@
 Notations: In what follows, $\|\cdot\|$ denotes the Euclidean norm on the relevant space, $\sigma_{\min}(M)$ is the smallest singular value of the matrix $M$.
-2 An adaptative switching algorithm
+2 An adapative switching algorithm
 We propose solving the smooth equality constrained problem
@@
 In (2.1), $f$ is a smooth function from $\hbox{I\thinspace R}^{n}$ into IR and $c$ is a smooth function from $\hbox{I\thinspace R}^{n}$ into $\hbox{I\thinspace R}^{m}$ ($m\leq n$).
-For a given vector $x$, we assume that we can compute the gradient $g(x)=\nabla_{x}f(x)$, as well as the (exact) Jacobian of the constraints at $x$
+For a given vector $x$, we assume that we can compute a random approximation $g(x)$ of the gradient $G(x)=\nabla_{x}f(x)$, as well ase the (exact) Jacobian of the constraints at $x$
 $J(x)=\nabla c(x)\in\hbox{I\thinspace R}^{m\times n}.$
-Assuming $J(x)$ has full rank, we may also compute the orthogonal projection onto its nullspace by
+and the orthogonal projection onto its nullspace by
 $P_{T}(x)\stackrel{{\scriptstyle\rm def}}{{=}}I-J(x)^{T}\big(J(x)J(x)^{T}\big)^{-1}J(x).$
-Given this projection, the projected gradient is then
-$g_{T}(x)\stackrel{{\scriptstyle\rm def}}{{=}}P_{T}(x)\,g(x)$ (2.2)
+Given this projection, the projected gradient and its approximation are then
+$G_{T}(x)\stackrel{{\scriptstyle\rm def}}{{=}}P_{T}(x)\,G(x)\;\;\mbox{ and }\;\;g_{T}(x)\stackrel{{\scriptstyle\rm def}}{{=}}P_{T}(x)\,g(x)$ (2.2)
 We may also define the (exact) least-squares Lagrange multiplier $\widehat{\lambda}(x)$ by
@@
 $L(x,\lambda)=f(x)+\lambda^{T}c(x)$ (2.4)
-and, for $\rho$ fixed, the associated “augmented-Lagrangian-like” Lyapunov function
+and, for $\rho$ fixed, the associated “augmented-Lagrangian-like” Lyapounov function
 $\psi_{\rho}(x,\lambda)\stackrel{{\scriptstyle\rm def}}{{=}}L(x,\lambda)+\rho\|c(x)\|,$ (2.5)
@@
 The definitions imply the important properties that
-$g_{T}(x)=\nabla f(x)+J(x)^{T}\widehat{\lambda}(x)\;\;\mbox{ and }\;\;J(x)\,g_{T}(x)=0,$ (2.7)
+$G_{T}(x)=\nabla f(x)+J(x)^{T}\widehat{\lambda}(x)\;\;\mbox{ and }\;\;J(x)\,g_{T}(x)=J(x)\,G_{T}(x)=0,$ (2.7)
 and also that, for $\lambda$ fixed to $\widehat{\lambda}(x)$,
@@
 By construction, we also have that
-$c(x)^{T}J(x)g_{T}(x)=0.$ (2.9)
+$c(x)^{T}J(x)g_{T}(x)=c(x)^{T}J(x)G_{T}(x)=0.$ (2.9)
 We are now in position to specify ADSWITCH in detail 2.
@@
 A starting point $x_{0}$ is given, together with constants $\omega>0$, $\theta>1$, $\beta,\eta,\varsigma\in(0,1]$.
-Set $\Gamma_{0}=0$ and $k=0$.
+Set $\Gamma_{-1}=0$ and $k=0$.
 Step 1: Evaluations.
-Evaluate $c_{k}=c(x_{k})$, $J_{k}=J(x_{k})$ and $g_{k}=\nabla_{x}f(x_{k})$.
+Evaluate $c_{k}=c(x_{k})$, $J_{k}=J(x_{k})$ and $g_{k}\approx\nabla_{x}f(x_{k})$.
 Set $g_{T,k}=P_{T}(x_{k})g_{k}$,
@@
 Step2: Tangential step: If $\|c_{k}\|\leq\beta\alpha_{T,k}\|g_{T,k}\|$ (2.12) then set $x_{k+1}=x_{k}-\alpha_{T,k}g_{T,k}$ (2.13) and $\Gamma_{k+1}=\Gamma_{k}^{+},$ (2.14) and go to Step 4.
-Step 3: Normal step: Otherwise, find a step $s_{N,k}$ in the range space of $J_{k}^{T}$ such that $\|s_{N,k}\|\leq\theta\|c_{k}\|$ (2.15) and there exists $\kappa_{\mbox{\tiny nrm}}>0$ independent of $k$ such that ${\scriptstyle\frac{1}{2}}\|c(x_{k}+s_{N,k})\|^{2}\leq{\scriptstyle\frac{1}{2}}\|c_{k}\|^{2}-\kappa_{\mbox{\tiny nrm}}\|J_{k}^{T}c_{k}\|^{2},$ (2.16) and set $x_{k+1}=x_{k}+s_{N,k}.$ (2.17) Step 4: Loop.
+Step 3: Normal step: Otherwise, find a step $s_{N,k}$ in the range space of $J_{k}^{T}$ such that $\|s_{N,k}\|\leq\theta\|J_{k}^{T}c_{k}\|$ (2.15) and there exists $\kappa_{\mbox{\tiny nrm}}>0$ independent of $k$ such that ${\scriptstyle\frac{1}{2}}\|c(x_{k}+s_{N,k})\|^{2}\leq{\scriptstyle\frac{1}{2}}\|c_{k}\|^{2}-\kappa_{\mbox{\tiny nrm}}\|J_{k}^{T}c_{k}\|^{2},$ (2.16) and set $x_{k+1}=x_{k}+s_{N,k}.$ (2.17) Step 4: Loop.
 Increment $k$ by one and go to Step 2.
   • The tangential step is the standard AdaGrad-norm [27] step in the nullspace of the Jacobian.
-  • The statement of the procedure to find the normal step in Step 3 may seem abstract because $\kappa_{\mbox{\tiny nrm}}$ may not be known, but, fortunately, there exist several standard techniques to achieve (2.15)-(2.16).
-    A first technique is to use a standard Armijo backtracking linesearch along the steepest-descent direction $J_{k}^{T}c_{k}$ (see Lemma 3.3).
-It is well-known that (2.16) holds in this case with $\kappa_{\mbox{\tiny nrm}}=1/L_{JTc}$ (see [3, Lemma 2.2.1], for instance).
+  • The statement of the procedure to find the normal step in Step 3 may seem abstract because $\kappa_{\mbox{\tiny nrm}}$ may not be known, but, fortunately, there exist several standard techniques to achieve (2.16).
+    A first technique is to use a standard Armijo backtracking linesearch along the the steepest-descent direction $J_{k}^{T}c_{k}$ with initial step size $\theta$.
+It is well-known that (2.16) holds in this case with $\kappa_{\mbox{\tiny nrm}}=1/L_{JTc}$ (see [2, Lemma 2.2.1], for instance).
 More generally, any “gradient related” direction $d_{k}$, that is any $d_{k}$ satisfying (2.15) and $d_{k}^{T}J_{k}^{T}c_{k}\leq-\kappa_{\mbox{\tiny grl}}\|J_{k}^{T}c_{k}\|^{2}$ for some $\kappa_{\mbox{\tiny grl}}>0$, can be used to initiate a backtracking linesearch.
-    A second possibility follows along this line and is to select a positive definite matrix $B_{k}$ with bounded condition number, to define $d_{k}=-B_{k}^{-1}J_{k}^{T}c_{k}$ and then to perform a backtracking linesearch along $d_{k}$ (see [5, Section 10.3.1] for a proof that $d_{k}$ is a gradient-related direction), in which case a similar result holds.
+    A second possibility follows along this line and is to select a positive definite matrix $B_{k}$ with bounded condition number, to define $d_{k}=-B_{k}^{-1}J_{k}^{T}c_{k}$ and then to perform a backtracking linesearch along $d_{k}$ (see [4, Section 10.3.1] for a proof that $d_{k}$ is a gradient-related direction), in which case a similar result holds.
 For instance, using $B_{k}=J_{k}^{T}J_{k}+\delta I_{n}$ for some $\delta\geq 0$ yields the regularized Gauss-Newton step
@@
 Moreover this is also the case for $\delta=0$ when $J_{k}$ is full-rank.)
-    Suitable variants of trust-region steps are also possible [5, Section 10.3.2], themselves involving several subvariants such as the truncated conjugate gradient [26, 27], CGLS [22] or LSQR [25].
+    Suitable variants of trust-region steps are also possible [4, Section 10.3.2], themselves involving several subvariants such as the truncated conjugate gradient [24, 25], CGLS [20] or LSQR [23].
 Because iterates generated by these methods are linear combinations of the gradients of ${\scriptstyle\frac{1}{2}}\|c_{k}+J_{k}s\|^{2}$, this guarantees that the final step belongs to the range of $J_{k}^{T}$.
@@
 In this case, if $Q_{k}$ and $R_{k}$ are the computed factors, then
-    $g_{T,k}=Q_{k,[1:n,m+1:n]}Q_{k,[1:n,m+1:n]}^{T}g_{k}$
+    $g_{N,k}=Q_{k[1;n,m+1:n]}Q_{k[1;n,m+1:n]}^{T}g_{k}$
     and the Newton step (2.18) is given by
-    $s_{N,k}=\left\{\begin{array}[]{ll}-Q_{k,[1:n,1:r]}R_{k,[1:r,1:r]}^{-T}c_{k}&\;\;\mbox{if }\;\;\delta=0,\\ -Q_{k,[1:n,1:m]}R_{k,[1:m,1:m]}(R_{k,[1:m,1:m]}^{T}R_{k,[1:m,1:m]}+\delta I_{m})^{-1}c_{k}&\;\;\mbox{otherwise.}\end{array}\right$.
+    $s_{N,k}=\left\{\begin{array}[]{ll}-Q_{k[1:n,1:r]}R_{k[1:r,1:r]}^{-T}c_{k}&\;\;\mbox{if }\;\;\delta=0,\\ -Q_{k[1:n,1:m}R_{k[1:m,1:m]}(R_{k[1:m,1:m]}^{T}R_{k[1:m,1:m]}+\delta I_{m})^{-1}c_{k}&\;\;\mbox{otherwise.}\end{array}\right$.
 (2.19)
-  • Observe that the ADSWITCH algorithm does not involve any explicit merit function or filter to control overall progress, but rely on the simple switching condition (2.12) (hence its name).
-As we will see below, an implicit merit function (a modified augmented-Lagrangian with suitably chosen parameters) is however crucial for our theoretical argument.
-  • ADSWITCH differs in several aspects from the trust-funnel methods of [17, 8].
+  • Observe that the ADSWITCH algorithm does not involve any explicit merit function or filter to control overal progress, but rely on the simple switching condition (2.12) (hence is name).
+As we will see below, an implicit merit function (an modified augmented-Lagrangian with suitably chosen parameters) is however crucial for our theoretical argument.
+  • ADSWITCH differs in several aspects from to the trust-funnel methods of [15, 7].
 The first is that it avoids evaluating the objective function in order to improve its reliability in the presence of noise.
@@
 Indeed it has been argued in [16] that the (unconstrained) AdaGrad algorithm can be viewed as a trust-region method with a specific radius management.
-Moreover, (2.15) can also be viewed as a trust-region constraint (such a condition is also imposed in [17]), but without the need to explicitly update its radius.
+Moreover, (2.15) can also be viewed as a trust-region constraint, but without the need to explicitly update its radius.
 Finally, ADSWITCH [^3] In its present incarnation.
-is more limited than trust-funnel methods in that it does not handle approximate projections in the tangent plane.
-In what follows, we denote by $\{k_{\tau}\}\subseteq\{k\}$ the index subsequence of “tangential iterations”, that is iterations where (2.12) holds.
-We also denote $\{k_{\nu}\}=\{k\}\setminus\{k_{\tau}\}$ the index subsequence of “normal iterations”.
+is more limited that trust-funnel methods in that it does not handle approximate projections in the tangent plane.
+In what follows, we denote by $\{k_{\tau}\}\subseteq\{k\}$ the index subsequence of tangential iterations, that is iterations where (2.12) holds.
+We alse denote $\{k_{\nu}\}=\{k\}\setminus\{k_{\tau}\}$ the index subsequence of normal iterations.
+We will use the abbreviations
+$\mathbb{E}^{\tau}_{k}\!\left[\cdot\right]\stackrel{{\scriptstyle\rm def}}{{=}}\mathbb{E}_{k}\!\left[\cdot|k\in\{k_{\tau}\}\right]\;\;\mbox{ and }\;\;\mathbb{E}^{\nu}_{k}\!\left[\cdot\right]\stackrel{{\scriptstyle\rm def}}{{=}}\mathbb{E}_{k}\!\left[\cdot|k\in\{k_{\nu}\}\right].$
 3 Complexity analysis
@@
 AS.1:
-For all $x\in\hbox{I\thinspace R}^{n}$, $f(x)\geq f_{\rm low}$.
+For all $k\geq 0$, $f(x_{k})\geq f_{\rm low}$.
 AS.2:
-For all $x\in\hbox{I\thinspace R}^{n}$, $\|g(x)\|\leq\kappa_{g}$ where $\kappa_{g}\geq\eta\beta$.
+For all $k\geq 0$, $\|G_{k}\|\leq\kappa_{g}$ where $\kappa_{g}\geq\eta^{2}\beta^{2}$.
 AS.3:
-For all $x\in\hbox{I\thinspace R}^{n}$, $\|c(x)\|\leq\kappa_{c}$.
+For all $k\geq 0$, $\|c_{k}\|\leq\kappa_{c}$.
 AS.4:
-For all $x\in\hbox{I\thinspace R}^{n}$, $\|J(x)\|\leq\kappa_{J}$
+For all $k\geq 0$, $\|J_{k}\|\leq\kappa_{J}$
 AS.5:
-For all $x\in\hbox{I\thinspace R}^{n}$, $\sigma_{\min}(J(x))\geq\sigma_{0}\in(0,1]$.
+For all $k\geq 0$, $\sigma_{\min}(J_{k})\geq\sigma_{0}\in(0,1]$.
 AS.6:
@@
 […]
-Note that these assumptions are very similar to those of [1, 7]. In particular, AS.5 is also part of Assumption 1 in this last reference.
-We finally assume, without loss of generality [^4] The parameter $\varsigma$ may be re-adjusted downwards at the first tangential iteration, if necessary., that
+Note that these assumptions are very similar to those of [6]. In particular, AS.5 is also part of Assumption 1 in this reference.
+Because of the random nature of the gradient estimator, the ADSWITCH algorithm generates a random process where, for a given iterate $x_{k}$, the oracle computes the gradient oracle $g_{k}=g(x_{k},\xi)$ where $\xi$ is a random variable (whose distribution may depend on $x_{k}$), with probability space $(\Omega,{\pazocal{F}},\mathbb{P})$.
+The expectation conditioned to knowing $g_{0},\ldots,g_{k-1}$ will be denoted by the symbol $\mathbb{E}_{k}\!\left[\cdot\right]$.
+Note that the projection $P_{T}(x_{k})$ is measurable with respect to the past, and so is $\widehat{\lambda}(x_{k})$.
+We complete our assumption as follows.
+AS.8 :
+For all $k\geq 0$, the estimator of $g_{k}$ is unbiased, that is $\mathbb{E}_{k}\!\left[g_{k}\right]=G_{k}$.
+AS.9 :
+For all $k\geq 0$, $\|g_{k}\|\leq\kappa_{g}$ where $\kappa_{g}\geq\eta^{2}\beta^{2}$.
+Observe that, because $P_{T}(x_{k})$ is a linear operator,
+$\mathbb{E}_{k}\!\left[g_{T,k}\right]=\mathbb{E}_{k}\!\left[P_{T}(x_{k})g_{k}\right]=P_{T}(x_{k})\mathbb{E}_{k}\!\left[g_{k}\right]=P_{T}(x_{k})G_{k}=G_{T,k}.$ (3.1)
+Using Jensen’s inequality and the convexity of the norm, we also have that
+$\mathbb{E}\!\left[\|G_{T,k}\|\right]=\mathbb{E}\!\left[\|\mathbb{E}_{k}\!\left[g_{T,k}\right]\|\right]\leq\mathbb{E}\!\left[\mathbb{E}_{k}\!\left[\|g_{T,k}\|\right]\right]=\mathbb{E}\!\left[\|g_{T,k}\|\right].$ (3.2)
+We next require a “directional root mean square” condition [1] along the tangential step $\alpha_{T,k_{\tau}}g_{T,k_{\tau}}$ given by
+AS.10:
+There exists a contant $\kappa_{\mbox{\tiny dir}}>0$ such that, for all $k\geq 0$,
+$\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}|(G_{T,k}-g_{T,k})^{T}g_{T,k}|\right]\leq\frac{\kappa_{\mbox{\tiny dir}}}{2}\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right].$
+We finally assume, without loss of generality [^4] The parameter $\varsigma$ may be re-adjusted downwards at the first tangential iteration, if necessary., that, for all realizations,
 $\Gamma_{k_{\tau}}\geq\varsigma\;\;\;\;(\tau\geq 0).$ (3.3)
 3.2 Tangential steps
-Our analysis hinges on the fact that first-order descent can be shown on the Lyapunov $\psi(x)$, both for tangential and normal steps, despite the fact that neither $\widehat{\lambda}(x_{k})$ or $\rho$ (which we still need to define) appears in the algorithm.
+Our analysis hinges on the fact that first-order descent can be shown on the Lyapounov $\psi(x)$, both for tangential and normal steps, despite the fact that neither $\widehat{\lambda}(x_{k})$ or $\rho$ (which we still need to define) appears in the algorithm.
 We start by considering tangential steps.
 Lemma 3.1
-Suppose that AS.4–AS.7 hold.
-Then $\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})\leq-\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|^{2}+\eta\,\kappa_{\mbox{\tiny tan}}\,\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$ (3.2) where $\kappa_{\mbox{\tiny tan}}=\frac{1}{2}\Big(L_{L}+\rho L_{c}+\beta L_{\lambda}+L_{\lambda}L_{c}\Big).$
+Suppose that AS.4–AS.10 hold.
+Then $\mathbb{E}^{\tau}_{k}\!\left[\psi(x_{k+1})\right]-\psi(x_{k})\leq-\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,kx}\|g_{T,k}\|^{2}\right]+\eta\,\kappa_{\mbox{\tiny tan}}\,\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right].$ (3.4) where $\kappa_{\mbox{\tiny tan}}=\frac{1}{2\eta}\Big(\kappa_{\mbox{\tiny dir}}+L_{L}+\rho L_{c}+\beta L_{\lambda}+L_{\lambda}L_{c}\Big).$
   • Proof.    We have that
-    $\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})$ $=\underbrace{\psi_{\rho}(x_{k_{\tau}+1},\widehat{\lambda}(x_{k_{\tau}}))-\psi_{\rho}(x_{k_{\tau}},\widehat{\lambda}(x_{k_{\tau}}))}_{\Delta_{x}}$ (3.3)
-    $\;\;\;\;\;\;\;\;\;\;\;\;+\underbrace{\psi_{\rho}(x_{k_{\tau}+1},\widehat{\lambda}(x_{k_{\tau}+1}))-\psi_{\rho}(x_{k_{\tau}+1},\widehat{\lambda}(x_{k_{\tau}}))}_{\Delta_{\lambda}}.$
+    $\mathbb{E}^{\tau}_{k}\!\left[\psi(x_{k+1})\right]-\psi(x_{k})$ $=\underbrace{\mathbb{E}^{\tau}_{k}\!\left[\psi_{\rho}(x_{k+1},\widehat{\lambda}(x_{k}))\right]-\psi_{\rho}(x_{k},\widehat{\lambda}(x_{k}))}_{\Delta_{x}}$ (3.5)
+    $\;\;\;\;\;\;\;\;\;\;\;\;+\underbrace{\mathbb{E}^{\tau}_{k}\!\left[\psi_{\rho}(x_{k+1},\widehat{\lambda}(x_{k+1}))-\psi_{\rho}(x_{k+1},\widehat{\lambda}(x_{k}))\right]}_{\Delta_{\lambda}}.$
 […]
-    $\Delta_{x}=-\nabla_{x}\psi_{\rho}(x_{k_{\tau}},\widehat{\lambda}(x_{k_{\tau}}))^{T}(\alpha_{T,k_{\tau}}g_{T,k_{\tau}})+r_{0}+\rho(\|c(x_{k_{\tau}+1})\|-\|c(x_{k_{\tau}})\|),$
+    $\Delta_{x}=-\mathbb{E}^{\tau}_{k}\!\left[(\nabla_{x}L_{\rho}(x_{k},\widehat{\lambda}(x_{k}))^{T}(\alpha_{T,k}g_{T,k})+r_{0}\right]+\rho(\mathbb{E}^{\tau}_{k}\!\left[\|c(x_{k+1})\|\right]-\|c(x_{k})\|),$
     with
-    $|r_{0}|\leq\frac{L_{L}}{2}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$
+    $\mathbb{E}^{\tau}_{k}\!\left[|r_{0}|\right]\leq\frac{L_{L}}{2}\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right].$
     Now
-    $\|c(x_{k_{\tau}+1})\|=\|c(x_{k_{\tau}})-\alpha_{T,k_{\tau}}J_{k}g_{T,k_{\tau}}+r_{1}\|\leq\|c(x_{k_{\tau}})\|+\|r_{1}\|$
+    $\mathbb{E}^{\tau}_{k}\!\left[\|c(x_{k+1})\|\right]=\mathbb{E}^{\tau}_{k}\!\left[\|c(x_{k})-\alpha_{T,k}J_{k}g_{T,k}+r_{1}\|\right]\leq\|c(x_{k})\|+\|r_{1}\|$
     with
-    $\|r_{1}\|\leq\frac{L_{c}}{2}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2},$
+    $\mathbb{E}^{\tau}_{k}\!\left[\|r_{1}\|\right]\leq\frac{L_{c}}{2}\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right],$
     giving
-    $\|c(x_{k_{\tau}+1})\|-\|c(x_{k_{\tau}})\|\leq\frac{L_{c}}{2}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$ (3.4)
-    Successively using (2.8) and (2.9), we obtain that
-    $\Delta_{x}$ $=-\alpha_{T,k_{\tau}}g_{T,k_{\tau}}^{T}g_{T,k_{\tau}}+r_{0}+\frac{\rho L_{c}}{2}\,\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}$ (3.5)
-    $\leq-\alpha_{T,k_{\tau}}g_{T,k_{\tau}}^{T}g_{T,k_{\tau}}+\frac{1}{2}\left(L_{L}+\rho L_{c}\right)\,\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}$
-    $\leq-\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|^{2}+\frac{1}{2}\left(L_{L}+\rho L_{c}\right)\,\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$
-    Now, we may use the Lipschitz continuity of $\widehat{\lambda}$ and $c$ to deduce that
-    $\Delta_{\lambda}$ $=c_{k_{\tau}+1}^{T}\big(\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\big)$ (3.6)
-    $=(c_{k_{\tau}+1}-c_{k_{\tau}})^{T}\big(\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\big)+c_{k_{\tau}}^{T}\big(\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\big)$
-    $\leq\|c_{k_{\tau}}\|\,\|\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\|+\|c_{k_{\tau}+1}-c_{k_{\tau}}\|\,\|\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\|$
-    $\leq L_{\lambda}\|c_{k_{\tau}}\|\,\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|+L_{\lambda}L_{c}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2},$
-    Taking (2.12) into account gives that $\|c_{k_{\tau}}\|\leq\beta\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|$, we obtain that
-    $\Delta_{\lambda}\leq\left(\beta L_{\lambda}+L_{\lambda}L_{c}\right)\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$
+    $\mathbb{E}^{\tau}_{k}\!\left[\|c(x_{k+1})\|\right]-\|c(x_{k})\|\leq\frac{L_{c}}{2}\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right].$ (3.6)
+    Successively using (2.8), (2.9), Jensens’ inequality and the convexity of the absolute value, and AS.10, we obtain that
+    $\Delta_{x}$ $=-\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}G_{T,k}^{T}g_{T,k}\right]+\mathbb{E}^{\tau}_{k}\!\left[r_{0}\right]+\frac{\rho L_{c}}{2}\,\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right]$ (3.7)
+    $\leq-\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}g_{T,k}^{T}g_{T,k}\right]+|\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}(G_{T,k}-g_{T,k})^{T}g_{T,k}\right]|+\frac{1}{2}\left(L_{l}+\rho L_{c}\right)\,\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right]$
+    $\leq-\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}g_{T,k}^{T}g_{T,k}\right]+\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}|G_{T,k}-g_{T,k})^{T}g_{T,k}|\right]+\frac{1}{2}\left(L_{l}+\rho L_{c}\right)\,\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right]$
+    $\leq-\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}\|g_{T,k}\|^{2}\right]+\frac{1}{2}\left(\kappa_{\mbox{\tiny dir}}+L_{L}+\rho L_{c}\right)\,\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right].$
+    Now, we may use the Lispchitz continuity of $\widehat{\lambda}$ and $c$ to deduce that
+    $\Delta_{\lambda}$ $=\mathbb{E}^{\tau}_{k}\!\left[c_{k+1}^{T}\big(\widehat{\lambda}(x_{k+1})-\widehat{\lambda}(x_{k})\big)\right]$ (3.8)
+    $=\mathbb{E}^{\tau}_{k}\!\left[(c_{k+1}-c_{k})^{T}\big(\widehat{\lambda}(x_{k+1})-\widehat{\lambda}(x_{k})\big)\right]+\mathbb{E}^{\tau}_{k}\!\left[c_{k}^{T}\big(\widehat{\lambda}(x_{k+1})-\widehat{\lambda}(x_{k})\big)\right]$
+    $\leq\mathbb{E}^{\tau}_{k}\!\left[\|c_{k}\|\,\|\widehat{\lambda}(x_{k+1})-\widehat{\lambda}(x_{k})\|\right]+\mathbb{E}^{\tau}_{k}\!\left[\|c_{k+1}-c_{k}\|\,\|\widehat{\lambda}(x_{k+1})-\widehat{\lambda}(x_{k})\|\right]$
+    $\leq L_{\lambda}\mathbb{E}^{\tau}_{k}\!\left[\|c_{k}\|\alpha_{T,k}\|g_{T,k}\|\right]+L_{\lambda}L_{c}\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right],$
+    Taking into account the fact that, for $k\in\{k_{\tau}\}$, (2.12) gives that $\|c_{k}\|\leq\beta\alpha_{T,k}\|g_{T,k}\|$, we obtain that
+    $\Delta_{\lambda}\leq\left(\beta L_{\lambda}+L_{\lambda}L_{c}\right)\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right].$
     Thus, summing $\Delta_{x}$ and $\Delta_{\lambda}$,
-    $\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})\leq$ $-\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|^{2}+\frac{1}{2}(L_{L}+\rho L_{c})\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}$
-    $+\left(\beta L_{\lambda}+L_{\lambda}L_{c}\right)\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}$
+    $\mathbb{E}^{\tau}_{k}\!\left[\psi(x_{k_{\tau}+1})\right]-\psi(x_{k_{\tau}})\leq$ $-\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|^{2}\right]+\frac{1}{2}(\kappa_{\mbox{\tiny dir}}+L_{L}+\rho L_{c})\mathbb{E}_{k}\!\left[\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}\right]$
+    $+\left(\beta L_{\lambda}+L_{\lambda}L_{c}\right)\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right]$
     and (3.4) follows. $\Box$
-The bound (3.2) quantifies the effect of tangential steps on the Lyapunov function, and its right-hand side involves a first-order (descent) term and a second-order perturbation term.
+The bound (3.4) quantifies the effect of tangential steps on the Lyapounov function, and its right-hand side involves a first-order (descent) term and a second-order perturbation term.
 We now derive crucial bounds on these terms.
@@
 Suppose that AS.2 and AS.5 hold.
-If we denote $\Gamma_{k_{\tau+1}}=\Gamma_{k_{\tau}}+\|g_{T,k_{\tau}}\|^{2}\;\;\mbox{ and }\;\;\alpha_{T,k_{\tau}}=\frac{\eta}{\sqrt{\Gamma_{k_{\tau+1}}+\varsigma}},$ then, for all $0\leq\tau_{0}\leq\tau_{1}$, $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}\,\|g_{T,k_{\tau}}\|^{2}>\frac{\eta}{2\sqrt{2}}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}$ (3.7) $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}^{2}\,\|g_{T,k_{\tau}}\|^{2}\leq\eta^{2}\,\log\left(\frac{\Gamma_{k_{\tau_{1}+1}}+\varsigma}{\Gamma_{k_{\tau_{0}}}+\varsigma}\right).$ (3.8)
-  • Proof.    Let $w_{k_{\tau}+1}=\sqrt{\Gamma_{k_{\tau+1}}+\varsigma}$. The definition (2.11) implies that
-    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}\,\|g_{T,k_{\tau}}\|^{2}$ $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{\|g_{T,k_{\tau}}\|^{2}}{\sqrt{\Gamma_{k_{\tau+1}}+\varsigma}}$
-    $>\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{\|g_{T,k_{\tau}}\|^{2}}{w_{k_{\tau+1}}+w_{k_{\tau}}}$
-    $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{w_{k_{\tau+1}}^{2}-w_{k_{\tau}}^{2}}{w_{k_{\tau+1}}+w_{k_{\tau}}}$
-    $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}(w_{k_{\tau+1}}-w_{k_{\tau}}),$
-    $=\eta(w_{k_{\tau_{1}+1}}-w_{k_{\tau_{0}}}),$
-    Now observe that, using (3.1),
-    $w_{k_{\tau_{1}+1}}-w_{k_{\tau_{0}}}$ $=\sqrt{\Gamma_{k_{\tau_{1}+1}}+\varsigma}-\sqrt{\Gamma_{k_{\tau_{0}}}+\varsigma}$
-    $=\frac{\Gamma_{k_{\tau_{1}+1}}-\Gamma_{k_{\tau_{0}}}}{\sqrt{\Gamma_{k_{\tau_{1}+1}}+\varsigma}+\sqrt{\Gamma_{k_{\tau_{0}}}+\varsigma}}$
-    $\geq\frac{\Gamma_{k_{\tau_{1}+1}}-\Gamma_{k_{\tau_{0}}}}{\sqrt{2\Gamma_{k_{\tau_{1}+1}}}+\sqrt{2\Gamma_{k_{\tau_{0}}}}}$
-    $>\frac{\Gamma_{k_{\tau_{1}+1}}}{2\sqrt{2\Gamma_{k_{\tau_{1}+1}}}}=\frac{\sqrt{\Gamma_{k_{\tau_{1}+1}}}}{2\sqrt{2}}$
-    which gives (3.7). Using the concavity and the increasing nature of the logarithm, we also have from (2.11) that
-    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}=\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\frac{\|g_{T,k_{\tau}}\|^{2}}{\Gamma_{k_{\tau+1}}+\varsigma}=\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\frac{\Gamma_{k_{\tau+1}}-\Gamma_{k_{\tau}}}{\Gamma_{k_{\tau+1}}+\varsigma}\leq\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\log(\Gamma_{k_{\tau+1}}+\varsigma)-\log(\Gamma_{k_{\tau}}+\varsigma),$
+If we denote $\Gamma_{k_{\tau+1}}=\Gamma_{k_{\tau}}+\|g_{T,k_{\tau}}\|^{2},\quad\alpha_{T,k_{\tau}}=\frac{\eta}{\sqrt{\Gamma_{k_{\tau}}+\varsigma}},\quad w_{k_{\tau}}=\sqrt{\Gamma_{k_{\tau}}+\varsigma^{2}},\quad$ then, for all $0\leq\tau_{0}\leq\tau_{1}$, $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}\,\|g_{T,k_{\tau}}\|^{2}>\eta\big(w_{k_{\tau_{1}}}-w_{k_{\tau_{0}-1}}\big),\$(3.9) $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}^{2}\,\|g_{T,k_{\tau}}\|^{2}\leq\eta^{2}\,\log\left(\frac{\Gamma_{k_{\tau_{1}}}+\varsigma}{\Gamma_{k_{\tau_{0}-1}}+\varsigma}\right),$ (3.10)
+  • Proof.    The definition (2.11) implies that, for every realization,
+    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}\,\|g_{T,k_{\tau}}\|^{2}$ $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{\|g_{T,k_{\tau}}\|^{2}}{\sqrt{\Gamma_{k_{\tau}}+\varsigma}}$
+    $>\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{\|g_{T,k_{\tau}}\|^{2}}{w_{k_{\tau}}+w_{k_{\tau}-1}}$
+    $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{w_{k_{\tau}}^{2}-w_{k_{\tau-1}}^{2}}{w_{k_{\tau}}+w_{k_{\tau-1}}}$
+    $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}(w_{k_{\tau}}-w_{k_{\tau-1}}),$
+    which gives (3.9) after taking the total expectation of both sides of the inequality. Using the concavity and the increasing nature of the logarithm, we also have from (2.11) that
+    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}=\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\frac{\|g_{T,k_{\tau}}\|^{2}}{\Gamma_{k_{\tau}}+\varsigma}=\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\frac{\Gamma_{k_{\tau}}-\Gamma_{k_{\tau-1}}}{\Gamma_{k_{\tau}}+\varsigma}\leq\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\log(\Gamma_{k_{\tau}}+\varsigma)-\log(\Gamma_{k_{\tau-1}}+\varsigma),$
     giving (3.10). $\Box$
 3.3 Normal steps
-Having considered tangential steps, we now show that a decrease in the Lyapunov function $\psi$ may also be proved if a normal step is taken, provided the penalty parameter $\rho$ is chosen large enough.
+Having considered tangential steps, we now show that a decrease in the Lyapounov function $\psi$ may also be proved if a normal step is taken, provided the penalty parameter $\rho$ is chosen large enough.
+Note that the normal step is deterministic.
 We start with a fairly simple observation.
 Lemma 3.3
-Suppose that AS.5 holds and that a normal step is used at iteration $k_{\nu}$. Then $\|c_{k_{\nu}+1}\|\leq\left(1-\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\right)\|c_{k_{\nu}}\|.$ (3.9)
+Suppose that AS.5 holds and that a normal step is used at iteration $k_{\nu}$. Then $\|c_{k_{\nu}+1}\|\leq\left(1-\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\right)\|c_{k_{\nu}}\|.$ (3.11)
   • Proof.    We have from (2.16) that $\|c_{k_{\nu}+1}\|<\|c_{k_{\nu}}\|$ Then
@@
     and therefore, using (2.16) again, that
-    $\|c_{k_{\nu}+1}\|-\|c_{k_{\nu}}\|\leq-\frac{\kappa_{\mbox{\tiny nrm}}\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|^{2}}{\|c_{k_{\nu}}\|}\leq-\frac{\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|^{2}}{\|c_{k_{\nu}}\|}=-\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|.$
+    $\|c_{k_{\nu}+1}\|-\|c_{k_{\nu}}\|\leq-\frac{\kappa_{\mbox{\tiny nrm}}\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|^{2}}{\|c_{k_{\nu}}\|}\leq-\frac{\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\|c_{k_{\nu}}\|^{2}}{\|c_{k_{\nu}}\|}=-\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\|c_{k_{\nu}}\|.$
     $\Box$
 Lemma 3.4
-Suppose that AS.3–AS.7 hold and that a normal step is used at iteration $k_{\nu}$.
-Define $\rho=\frac{1}{\kappa_{\rm nrm}\,\sigma_{0}^{2}}\Big[\;\kappa_{c}\,\theta\,\,L_{\lambda}\;+\;\frac{1}{2}\,\kappa_{c}\,\theta^{2}\,(L_{L}+L_{\lambda}L_{c})\;+\;\eta\;\Big]$ (3.10) Then we have that $\psi(x_{k_{\nu}+1})-\psi(x_{k_{\nu}})\leq-\eta\|c_{k_{\nu}}\|.$ (3.11)
+Suppose that AS.4–AS.7 hold and that a normal step is used at iteration $k_{\nu}$.
+Define $\rho=\frac{\kappa_{c}\theta}{\kappa_{\mbox{\tiny nrm}}\kappa_{J}}\left[L_{\lambda}+\frac{\theta}{2}(L_{L}+L_{\lambda}L_{c})+\eta\right].$ (3.12) Then we have that $\psi(x_{k_{\nu}+1})-\psi(x_{k_{\nu}})\leq-\eta\|c_{k_{\nu}}\|.$ (3.13)
 […]
@@
     $=L(x_{k_{\nu}+1},\widehat{\lambda}_{k_{\nu}})-L(x_{k_{\nu}},\widehat{\lambda}_{k_{\nu}})+\rho\big(\|c_{k_{\nu}+1}\|-\|c_{k_{\nu}}\|\big)$
-    $\leq(\nabla_{x}L(x_{k_{\nu}},\widehat{\lambda}_{k_{\nu}})^{T}s_{N,k_{\nu}}+r_{3}-\rho\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|$
-    $\leq-\rho\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|+\frac{\theta^{2}L_{L}}{2}\|c_{k_{\nu}}\|^{2}.$
+    $\leq(\nabla_{x}L(x_{k_{\nu}},\widehat{\lambda}_{k_{\nu}})^{T}s_{N,k_{\nu}}+r_{3}-\rho\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\|c_{k_{\nu}}\|$
+    $\leq-\rho\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\|c_{k_{\nu}}\|+\frac{\theta^{2}L_{L}}{2}\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|^{2}.$
     Using AS.6, the Lipschitz continuity of $\widehat{\lambda}$ and $c$ and AS.5, we then deduce that
-    $\Delta_{\lambda}$ $=\psi_{\rho}(x_{k_{\nu}+1},\widehat{\lambda}_{k_{\nu}+1})-\psi_{\rho}(x_{k_{\nu}+1},\widehat{\lambda}_{k_{\nu}})$ (3.13)
+    $\Delta_{\lambda}$ $=\psi_{\rho}(x_{k_{\nu}+1},\widehat{\lambda}(x_{k_{\nu}+1}))-\psi_{\rho}(x_{k_{\nu}+1},\widehat{\lambda}(x_{k_{\nu}}))$ (3.15)
     $\leq(\|c_{k_{\nu}}\|+\|c_{k_{\nu}+1}-c_{k_{\nu}}\|)\,\|\widehat{\lambda}_{k_{\nu}+1}-\widehat{\lambda}_{k_{\nu}}\|$
     $\leq L_{\lambda}\,\|s_{N,k_{\nu}}\|\,\|c_{k_{\nu}}\|+\frac{L_{\lambda}L_{c}}{2}\|s_{N,k_{\nu}}\|^{2}$
-    $\leq L_{\lambda}\,\theta\|c_{k_{\nu}}\|^{2}+\frac{L_{\lambda}L_{c}\theta^{2}}{2}\|c_{k_{\nu}}\|^{2}$
+    $\leq L_{\lambda}\,\theta\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|\,\|c_{k_{\nu}}\|+\frac{L_{\lambda}L_{c}\theta^{2}}{2}\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|^{2}$
     and thus, summing (3.14) and (3.15), that
-    $\psi(x_{k_{\nu}+1})-\psi(x_{k_{\nu}})\leq-\rho\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|+L_{\lambda}\,\theta\kappa_{c}\,\|c_{k_{\nu}}\|+\frac{1}{2}\left(L_{L}+L_{\lambda}L_{c}\right)\theta^{2}\kappa_{c}\|c_{k_{\nu}}\|.$
-    The bound (3.11) then follows from (3.10).
-    $\Box$
+    $\psi(x_{k_{\nu}+1})-\psi(x_{k_{\nu}})$ $\leq-\rho\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\|c_{k_{\nu}}\|+L_{\lambda}\,\theta\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|\,\|c_{k_{\nu}}\|+\left(\frac{\theta^{2}L_{L}}{2}+\frac{L_{\lambda}L_{c}\theta^{2}}{2}\right)\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|^{2}$
+    $\leq-\rho\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\|c_{k_{\nu}}\|+L_{\lambda}\,\theta\kappa_{J}\kappa_{c}\,\|c_{k_{\nu}}\|+\left(\frac{\theta^{2}L_{L}}{2}+\frac{L_{\lambda}L_{c}\theta^{2}}{2}\right)\kappa_{J}^{2}\kappa_{c}\|c_{k_{\nu}}\|.$
+    The bound (3.13) then follows from (3.12). $\Box$
 3.4 Telescoping sum
 We now combine the effects of tangential and normal steps to accumulate in a telescoping sum the relevant decreases on $\psi(x)$ across all iterations.
+This requires an additional assumption involving where
+$\mathbb{P}_{T,j}\stackrel{{\scriptstyle\rm def}}{{=}}\mathbb{P}_{j}\!\left[\frac{\alpha_{T,j}\|g_{j}\|}{\|c_{j}\|}\right]$
+(We assume that $\mathbb{P}_{T,j}=1$ if $\|c_{j}\|=0$.) Note that $\mathbb{P}_{T,j}$ is the conditional probability that iteration $j$ uses a tangential step, i.e.
+the conditional probability that $j\in\{k_{\tau}\}$.
+We assume that
+AS.11:
+There exists a constant $\mu_{T}\in(0,1]$ such that
+$\sum_{j}\mathbb{E}_{j}\!\left[\alpha_{T,j}\|g_{T,j}\|^{2}\mathbb{P}_{T,j}\right]\geq\mu_{T}\sum_{j\in{k_{\tau}}}\mathbb{E}_{j}\!\left[\alpha_{T,j}\|g_{T,j}\|^{2}\right]$
 Lemma 3.5
-Suppose that AS.0–AS.7 hold.
-Then $\sqrt{\Gamma_{k_{\tau_{1}+1}}}+\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right),$ (3.14) where $\kappa_{\mbox{\tiny gap}}=\frac{2\sqrt{2}}{\eta}\left(\psi(x_{0})+\frac{\kappa_{J}\kappa_{g}\kappa_{c}}{\sigma_{0}^{2}}+\rho\kappa_{c}-f_{\rm low}\right).$
-  • Proof.    Combining (3.2), (3.7) and (3.8), we obtain that
-    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\left(\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})\right)$ $\leq-\frac{\eta}{2\sqrt{2}}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}+\eta^{2}\kappa_{\mbox{\tiny tan}}\log\left(\frac{\Gamma_{k_{\tau_{1}+1}}+\varsigma}{\Gamma_{k_{\tau_{0}}}+\varsigma}\right)$ (3.15)
-    Then, if $\min[k_{\nu_{0}},k_{\tau_{0}}]=0$ and $\max[k_{\nu_{1}},k_{\tau_{1}}]=k$, we have that $w_{k_{\tau_{0}}}=\varsigma$ and $\Gamma_{k_{\tau_{0}}}=0$.
+Suppose that AS.0–AS.10 hold.
+Then $\mathbb{E}\!\left[\sqrt{\Gamma_{k_{\tau_{1}}}}\right]+\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(1+\frac{\Gamma_{k_{\tau_{1}}}}{\varsigma}\right)\right],$ (3.16) where $\kappa_{\mbox{\tiny gap}}=\frac{1}{\eta}\left(\psi(x_{0})+\frac{\kappa_{J}\kappa_{g}\kappa_{c}}{\sigma_{0}}-f_{\rm low}\right)+\varsigma.$
+  • Proof.    Consider first the effect of tangential steps. We have that
+    $\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\tau}_{j}\!\left[\left(\psi(x_{j+1})-\psi(x_{j})\right)\right]\mathbb{P}_{T,j}\right]$ $\leq\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\tau}_{j}\!\left[\left(-\alpha_{T,j}\|g_{T,j}\|^{2}+\alpha_{T,j}^{2}\|g_{T,j}\|^{2}\right)\right]\mathbb{P}_{T,j}\right]$
+    $=\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\left(-\alpha_{T,j}\|g_{T,j}\|^{2}+\alpha_{T,j}^{2}\|g_{T,j}\|^{2}\right)\mathbb{P}_{T,j}\right]$
+    where we used (3.4) to derive the first inequality and the independence of the two gradient estimations to deduce the final equality. Now AS.11, (3.9) and (3.10) give that
+    $\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\tau}_{j}\!\left[\left(\psi(x_{j+1})-\psi(x_{j})\right)\right]\mathbb{P}_{T,j}\right]$ $\leq=-\mu_{T}\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\alpha_{T,j}\|g_{T,j}\|^{2}\right]+\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\alpha_{T,j}^{2}\|g_{T,j}\|^{2}\right]$
+    $\leq\eta\mu_{T}\,\mathbb{E}_{k}\!\left[w_{k_{\tau_{1}}}-w_{k_{\tau_{0}-1}}\right]+\eta\mu_{T}\kappa_{\mbox{\tiny tan}}\mathbb{E}_{j}\!\left[\log\left(\frac{\Gamma_{k_{\tau_{1}}}+\varsigma}{\Gamma_{k_{\tau_{0}-1}}+\varsigma}\right)\right]$
+    For normal steps,
+    $\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\nu}_{j}\!\left[\left(\psi(x_{j+1})-\psi(x_{j})\right)\right](1-\mathbb{P}_{T,j})\right]=-\eta\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\|c_{j}\|(1-\mathbb{P}_{T,j})\right]\leq-\eta\mu_{N}\sum_{j=0}^{k}\|c_{j}\|\mathbb{E}_{j}\!\left[(1-\mathbb{P}_{T,j})\right]$
+    Combining (3.4), (3.9) and (3.10) and taking total expectation, we obtain that
+    $\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\left(\psi(x_{j+1})\right]-\psi(x_{j})\right)$ $=\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\tau}_{j}\!\left[\left(\psi(x_{j+1})-\psi(x_{j})\right)\right]\mathbb{P}_{T,j}\right]$ (3.17)
+    $=\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\tau}_{j}\!\left[\left(\psi(x_{j+1})-\psi(x_{j})\right)\right]\right]\mathbb{E}_{j}\!\left[\mathbb{P}_{T,j}\right]$
+    $\hskip 28.45274pt+\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\nu}_{j}\!\left[\left(\psi(x_{j+1})-\psi(x_{j})\right)\right](1-\mathbb{P}_{T,j})\right]$
+    $\leq\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\left(-\alpha_{T,j}|g_{T,j}\|^{2}+\alpha_{T,j}^{2}\|g_{T,j}\|^{2}\right)\right]\mathbb{E}_{j}\!\left[\mathbb{P}_{T,j}\right]$
+    $\hskip 28.45274pt+\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[(-\eta\|c_{j}\|)(1-\mathbb{P}_{T,j})\right]$
+    $\leq\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[-\alpha_{T,j}|g_{T,j}\|^{2}\right]\mathbb{E}_{j}\!\left[\mathbb{P}_{T,j}\right]+\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\alpha_{T,j}^{2}\|g_{T,j}\|^{2}\right]$
+    $\hskip 28.45274pt+\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[(-\eta\|c_{j}\|)(1-\mathbb{P}_{T,j})\right]$
+    $\leq\mu_{T}\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[-\alpha_{T,j}|g_{T,j}\|^{2}\right]+\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\alpha_{T,j}^{2}\|g_{T,j}\|^{2}\right]$
+    $\hskip 28.45274pt+\mu_{N}\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[(-\eta\|c_{j}\|)\right]$
+    $\leq-\eta\mu\,\mathbb{E}\!\left[w_{k_{\tau_{1}}}-w_{k_{\tau_{0}-1}}\right]+\eta\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(\frac{\Gamma_{k_{\tau_{1}}}+\varsigma}{\Gamma_{k_{\tau_{0}-1}}+\varsigma}\right)\right]+???$
+    Then, if $\min[k_{\nu_{0}},k_{\tau_{0}}]=0$ and $\max[k_{\nu_{1}},k_{\tau_{1}}]=k$, we have that $w_{k_{\tau_{0}-1}}=\varsigma$ and $\Gamma_{k_{\tau_{0}-1}}=0$.
 The bounds (3.13) and (3.17) therefore give that
-    $\psi(x_{k+1})$ $-\psi(x_{0})$ (3.16)
-    $=\sum_{\tau=\tau_{0}}^{\tau_{1}}\big(\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})\big)+\sum_{\nu=\nu_{0}}^{\nu_{1}}\big(\psi(x_{k_{\nu}+1})-\psi(x_{k_{\nu}})\big)$
-    $\leq-\frac{\eta}{2\sqrt{2}}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}-\eta\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|+\eta^{2}\kappa_{\mbox{\tiny tan}}\log\left(\frac{\Gamma_{k_{\tau_{1}+1}}+\varsigma}{\Gamma_{k_{\tau_{0}}}+\varsigma}\right)$
-    $\leq-\frac{\eta}{2\sqrt{2}}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}-\frac{\eta}{2\sqrt{2}}\,\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|+\eta\kappa_{\mbox{\tiny tan}}\log\left(1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right)$
-    where we used the fact that $\eta\leq 1$. But (2.3), AS.2, AS.4, and AS.5 ensure that
-    $\|\widehat{\lambda}(x)\|\leq\frac{\kappa_{J}}{\sigma_{0}^{2}}\|g(x)\|\leq\frac{\kappa_{J}\kappa_{g}}{\sigma_{0}^{2}}.$
-    Hence, using (2.5), AS.1, AS.2, AS.3, we have that
-    $\psi(x_{k+1})-\psi(x_{0})\geq f_{\rm low}-\frac{\kappa_{J}\kappa_{g}\kappa_{c}}{\sigma_{0}^{2}}-\rho\kappa_{c}-\psi(x_{0})\stackrel{{\scriptstyle\rm def}}{{=}}-\eta\kappa_{\mbox{\tiny gap}},$
+    $\mathbb{E}\!\left[\psi(x_{k+1})\right]$ $-\psi(x_{0})$ (3.18)
+    $=\sum_{\tau=\tau_{0}}^{\tau_{1}}\big(\mathbb{E}\!\left[\psi(x_{k_{\tau}+1})\right]-\mathbb{E}\!\left[\psi(x_{k_{\tau}})\right]\big)+\sum_{\nu=\nu_{0}}^{\nu_{1}}\big(\mathbb{E}\!\left[\psi(x_{k_{\nu}+1})\right]-\mathbb{E}\!\left[\psi(x_{k_{\nu}})\right]\big)$
+    $\leq-\eta\,\mathbb{E}\!\left[w_{k_{\tau_{1}}}-w_{k_{\tau_{0}-1}}\right]-\eta\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]+\eta\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(\frac{\Gamma_{k_{\tau_{1}}}+\varsigma}{\Gamma_{k_{\tau_{0}-1}}+\varsigma}\right)\right]$
+    $\leq-\eta\left(\mathbb{E}\!\left[\sqrt{\Gamma_{k_{\tau_{1}}}}\right]-\sqrt{\varsigma}\right)-\eta\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]+\eta\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(1+\frac{\Gamma_{k_{\tau_{1}}}}{\varsigma}\right)\right]$
+    where we used AS.5. But AS.4, AS.5 and AS.7 ensure that
+    $\widehat{\lambda}(x)\leq\frac{\kappa_{J}}{\sigma_{0}}\|G(x)\|\leq\frac{\kappa_{J}\kappa_{g}}{\sigma_{0}}.$
+    Hence, using (2.5), AS.1, AS.2, AS.3, we have that, for all realizations,
+    $\psi(x_{k+1})-\psi(x_{0})\geq f_{\rm low}-\frac{\kappa_{J}\kappa_{g}\kappa_{c}}{\sigma_{0}}-\psi(x_{0})\stackrel{{\scriptstyle\rm def}}{{=}}-\eta(\kappa_{\mbox{\tiny gap}}+\varsigma),$
     so that (3.18) implies (3.16). $\Box$
-This result is central in our analysis as it provides consistent upper bounds on a quantity related to the optimality mesure $\|g_{T,k}\|+\|c_{k}\|$, which we now exploit separately.
+This result is central in our analysis as it provides consistent upper bounds on a quantity related to the expected value of the optimality mesure $\|g_{T,k}\|+\|c_{k}\|$, which we now exploit separately.
 3.5 Tangential complexity
@@
 Suppose that $at\leq b+c\,\log(t)$ for $t\geq 1$. Then $t\leq\max\left[e^{b/c},\frac{4c^{2}}{a^{2}}\right].$
-  • Proof.    Suppose that $t>e^{b/c}$. Then $b\leq c\,\log(t)$ and thus (see [28, eq. (14)])
+  • Proof.    Suppose that $t>e^{b/c}$. Then $b\leq c\,\log(t)$ and thus (see [26, eq. (14)])
     $at\leq 2c\,\log(t)\leq 2c\,\log(1+t)\leq\frac{2ct}{\sqrt{1+t}}.$
@@
 Lemma 3.7
-Suppose that AS.0–AS.7 hold.
-Then $\sqrt{\Gamma_{k_{\tau_{1}+1}}}\leq\kappa_{T}\stackrel{{\scriptstyle\rm def}}{{=}}\sqrt{\frac{\varsigma}{2}}\max\left[e^{\frac{\kappa_{\mbox{\tiny gap}}}{2\kappa_{\mbox{\tiny tan}}}},\frac{32\kappa_{\mbox{\tiny tan}}^{2}}{\varsigma}\right]$ (3.17) and $\sum_{\tau=\tau_{0}}^{\tau_{1}}\|g_{T,k_{\tau}}\|+\|c_{k_{\tau}}\|\leq\kappa_{T}\sqrt{\tau_{1}+1}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right).$ (3.18)
+Suppose that AS.0–AS.10 hold.
+Then $\mathbb{E}\!\left[\sqrt{\Gamma_{k_{\tau_{1}}}}\right]\leq\kappa_{T}\stackrel{{\scriptstyle\rm def}}{{=}}\sqrt{\frac{\varsigma}{2}}\max\left[e^{\frac{\kappa_{\mbox{\tiny gap}}}{\kappa_{\mbox{\tiny tan}}}},2\kappa_{\mbox{\tiny gap}}^{2}\right]$ (3.19) and $\sum_{\tau=\tau_{0}}^{\tau_{1}}\mathbb{E}\!\left[\|G_{T,k_{\tau}}\|+\|c_{k_{\tau}}\|\right]\leq\kappa_{T}\sqrt{\tau_{1}+1}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right)$ (3.20)
   • Proof.    The bound (3.16) implies that
-    $\sqrt{\Gamma_{k_{\tau_{1}+1}}}\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right).$
+    $\mathbb{E}\!\left[\sqrt{\Gamma_{k_{\tau_{1}}}}\right]\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(1+\frac{\Gamma_{k_{\tau_{1}}}}{\varsigma}\right)\right].$
     Now (3.3) implies that
-    $1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\leq\frac{2}{\varsigma}\Gamma_{k_{\tau_{1}+1}}$
+    $1+\frac{\Gamma_{k_{\tau_{1}}}}{\varsigma}\leq\frac{2}{\varsigma}\Gamma_{k_{\tau_{1}}}$
     and thus that
-    $\sqrt{\frac{\varsigma}{2}}\sqrt{\frac{2\Gamma_{k_{\tau_{1}+1}}}{\varsigma}}\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(\frac{2\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right)=\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\sqrt{\frac{2\Gamma_{k_{\tau_{1}+1}}}{\varsigma}}\right).$
+    $\sqrt{\frac{\varsigma}{2}}\mathbb{E}\!\left[\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right]\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}\right)\right]=\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right)\right].$
+    We may now use Jensen’s inequality and the concavity of the logarithm to deduce that
+    $\sqrt{\frac{\varsigma}{2}}\mathbb{E}\!\left[\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right]\leq\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\mathbb{E}\!\left[\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right]\right).$
     Using Lemma 3.6, we then obtain that
-    $\sqrt{\Gamma_{k_{\tau_{1}+1}}}=\sqrt{\frac{\varsigma}{2}}\sqrt{\frac{2\Gamma_{k_{\tau_{1}+1}}}{\varsigma}}\leq\sqrt{\frac{\varsigma}{2}}\max\left[e^{\frac{\kappa_{\mbox{\tiny gap}}}{2\kappa_{\mbox{\tiny tan}}}},\frac{32\kappa_{\mbox{\tiny tan}}^{2}}{\varsigma}\right]$
+    $\mathbb{E}\!\left[\sqrt{\Gamma_{k_{\tau_{1}}}}\right]=\sqrt{\frac{\varsigma}{2}}\mathbb{E}\!\left[\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right]\leq\sqrt{\frac{\varsigma}{2}}\max\left[e^{\frac{\kappa_{\mbox{\tiny gap}}}{\kappa_{\mbox{\tiny tan}}}},2\kappa_{\mbox{\tiny gap}}^{2}\right]$
     This is (3.19). We may now invoke the inequality
@@
     for nonnegative $\{a_{j}\}_{j=0}^{k}$ to deduce from (2.10) and (3.19) that
-    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\|g_{T,k_{\tau}}\|\leq\sqrt{\tau_{1}+1}\,\sqrt{\sum_{\tau=\tau_{0}}^{\tau_{1}}\|g_{T,k_{\tau}}\|^{2}}=\sqrt{\tau_{1}+1}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}\leq\sqrt{\tau_{1}+1}\,\kappa_{T}.$ (3.19)
+    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\mathbb{E}\!\left[\|g_{T,k_{\tau}}\|\right]\leq\sqrt{\tau_{1}+1}\,\mathbb{E}\!\left[\sqrt{\sum_{j=0}^{k}\|g_{T,k_{\tau}}\|^{2}}\right]=\sqrt{\tau_{1}+1}\,\mathbb{E}\!\left[\sqrt{\Gamma_{k_{\tau_{1}}}}\right]\leq\sqrt{\tau_{1}+1}\,\kappa_{T},$
+    which, with (3.2), gives that
+    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\mathbb{E}\!\left[\|G_{T,k_{\tau}}\|\right]\leq\kappa_{T}\sqrt{\tau_{1}+1}.$ (3.21)
     Using the switching condition (2.12) and the fact that $\alpha_{T,k_{\tau}}\leq\eta/\sqrt{\varsigma}$, we also deduce that
-    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\|c_{k_{\tau}}\|\leq\sum_{\tau=\tau_{0}}^{\tau_{1}}\beta\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|\leq\frac{\beta\eta}{\sqrt{\varsigma}}\sum_{\tau=\tau_{0}}^{\tau_{1}}\|g_{T,k_{\tau}}\|\leq\frac{\beta\eta\kappa_{T}\sqrt{\tau_{1}+1}}{\sqrt{\varsigma}}.$
+    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\mathbb{E}\!\left[\|c_{k_{\tau}}\|\right]\leq\sum_{\tau=\tau_{0}}^{\tau_{1}}\beta\mathbb{E}\!\left[\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|\right]\leq\frac{\beta\eta}{\sqrt{\varsigma}}\sum_{\tau=\tau_{0}}^{\tau_{1}}\mathbb{E}\!\left[\|g_{T,k_{\tau}}\|\right]\leq\frac{\beta\eta\kappa_{T}\sqrt{\tau_{1}+1}}{\sqrt{\varsigma}}.$
     Summing this bound with (3.21) then gives (3.20). $\Box$
-Note that a marginally tighter bound on $\Gamma_{k_{\tau_{1}}}$ can obtained, as in [2], by using the Lambert $W_{-1}$ function (see [6]) instead of applying Lemma 3.6, but we have chosen the latter for simplicity of exposition.
+Note that a marginally tighter bound on $\mathbb{E}\!\left[\Gamma_{k_{\tau_{1}}}\right]$ can obtained, as in [1], by using the Lambert $W_{-1}$ function (see [5]) instead of applying Lemma 3.6, but we have chosen the latter for simplicity of exposition.
 3.6 Normal complexity
@@
 Lemma 3.8
-Suppose that AS.0–AS.7 hold.
-Then $\sum_{\nu=\nu_{0}}^{\nu_{1}}\|g_{T,k_{\nu}}\|+\|c_{k_{\nu}}\|<\kappa_{N}\,\left(1+\frac{2\sqrt{2}}{\beta\eta}\right),$ (3.20) where $\kappa_{N}=\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(\frac{2\kappa_{T}}{\varsigma}\right).$ (3.21)
+Suppose that AS.0–AS.10 hold.
+Then $\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|G_{T,k_{\nu}}\|+\|c_{k_{\nu}}\|\right]<\frac{4\kappa_{N}\,\kappa_{g}}{\beta\eta}\sqrt{k_{\nu}+1},$ (3.22) where $\kappa_{N}=\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(\frac{2\kappa_{T}}{\varsigma}\right).$ (3.23)
   • Proof.    The bound (3.16) ensures that
-    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right)$ (3.22)
-    where $k_{\tau_{1}}$ is the index of the last tangential iteration before $k_{\nu_{1}}$.
-As in Lemma 3.7, we may now use the assumption that $\Gamma_{k_{\tau_{1}}}\geq\varsigma$ and the monotonicity of the logarithm to obtain that
-    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|\leq\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right)\leq\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\sqrt{\frac{2}{\varsigma}}\kappa_{T}\right)=\kappa_{N}.$ (3.23)
-    Using the switching condition (2.12), we obtain that, for $\nu\in\{\nu_{0},\ldots,\nu_{1}\}$,
-    $\|c_{k_{\nu}}\|\geq\beta\,\alpha_{T,k_{\nu}}\|g_{T,k_{\nu}}\|.$ (3.24)
-    Observe now that, using the assumption that $\Gamma_{k_{\nu}}\geq\varsigma$, and (3.17),
-    $\alpha_{T,k_{\nu}}=\frac{\eta}{\sqrt{\Gamma_{k_{\nu}}+\varsigma}}\geq\frac{\eta}{\sqrt{2\Gamma_{k_{\nu}}}}\geq\frac{\eta}{\sqrt{2\kappa_{T}}}.$
-    Thus, (3.24) yields that, for $\nu\in\{\nu_{0},\ldots,\nu_{1}\}$
-    $\|c_{k_{\nu}}\|\ \geq\frac{\beta\eta}{\sqrt{2\kappa_{T}}}\|g_{T,k_{\nu}}\|.$ (3.25)
+    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(1+\frac{\Gamma_{k_{\tau_{1}}}}{\varsigma}\right)\right]$ (3.24)
+    where $k_{\tau_{1}}$ is the index of the last tangental iteration before $k_{\nu_{1}}$.
+As in Lemma 3.7, we may now use the assumption that $\Gamma_{k_{\tau_{1}}}\geq\varsigma$, Jensen’s inequality and the concavity and monotonicity of the logarithm to obtain that
+    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]\leq\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\mathbb{E}\!\left[\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right]\right)\leq\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\sqrt{\frac{2}{\varsigma}}\kappa_{T}\right)=\kappa_{N}.$ (3.25)
+    Taking now the conditional expectation in the switching condition (2.12), we obtain that, for $\nu\in\{\nu_{0},\ldots,\nu_{1}\}$,
+    $\|c_{k_{\nu}}\|\geq\beta\,\mathbb{E}_{k_{\nu}}\!\left[\alpha_{T,k_{\nu}}\|g_{T,k_{\nu}}\|\right].$
+    Observe now that, using the assumption that $\Gamma_{k_{\nu}}\geq\varsigma$, (2.10) and AS.9,
+    $\alpha_{T,k_{\nu}}=\frac{\eta}{\sqrt{\Gamma_{k_{\nu}}+\varsigma}}\geq\frac{\eta}{\sqrt{2\Gamma_{k_{\nu}}}}\geq\frac{\eta}{\kappa_{g}\sqrt{2(k_{\nu}+1)}}.$
+    Thus, using (3.2),
+    $\|c_{k_{\nu}}\|\ \geq\frac{\beta\eta}{\kappa_{g}\sqrt{2(k_{\nu}+1)}}\mathbb{E}_{k_{\nu}}\!\left[\|g_{T,k_{\nu}}\|\right]\geq\frac{\beta\eta}{\kappa_{g}\sqrt{2(k_{\nu}+1)}}\|G_{T,k_{\nu}}\|,$
+    in turn yielding that
+    $\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]\geq\frac{\beta\eta}{\kappa_{g}\sqrt{2(k_{\nu}+1)}}\mathbb{E}\!\left[\|G_{T,k_{\nu}}\|\right].$ (3.26)
     With (3.25), this implies that
-    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\|g_{T,k_{\nu}}\|\leq\frac{\sqrt{2\kappa_{T}}}{\beta\eta}\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|\leq\frac{\kappa_{N}\,\sqrt{2\kappa_{T}}}{\beta\eta}.$
-    Summing this bound with (3.23) gives (3.20). $\Box$
+    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|G_{k_{\nu}}\|\right]\leq\frac{\kappa_{g}\sqrt{2(k_{\nu}+1)}}{\beta\eta}\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]\leq\frac{\kappa_{N}\,\kappa_{g}\sqrt{2(k_{\nu}+1)}}{\beta\eta}.$
+    Summing this bound with (3.25) and using $\sqrt{\kappa_{g}}\leq\beta\eta$ gives (3.22). $\Box$
 3.7 Combined complexity
@@
 Theorem 3.9
-Suppose that AS.0-AS.7 hold.
-Then $\frac{1}{k+1}\sum_{j=0}^{k}\Big(\|g_{T,j}\|+\|c_{j}\|\Big)\leq\frac{\kappa_{\mbox{\tiny ADSW,1}}}{\sqrt{k+1}}+\frac{\kappa_{\mbox{\tiny ADSW,2}}}{k+1}={\pazocal{O}}\left(\frac{1}{\sqrt{k+1}}\right),$ (3.26) where $\kappa_{\mbox{\tiny ADSW,1}}=\kappa_{T}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right)+\frac{4\kappa_{N}\,\kappa_{g}}{\beta\eta}\;\;\mbox{ and }\;\;\kappa_{\mbox{\tiny ADSW,2}}=\kappa_{N}\,\left(1+\frac{2\sqrt{2}}{\beta\eta}\right).$
+Suppose that AS0-AS7 hold and that either the gradient is exact (i.e.
+$g_{T,k}=G_{T,k}$ for all $k\geq 0$) or AS.8–AS.10 hold.
+Then $\frac{1}{k+1}\sum_{j=0}^{k}\mathbb{E}\!\left[\|G_{T,j}\|+\|c_{j}\|\right]\leq\frac{\kappa_{\mbox{\tiny ADSW}}}{\sqrt{k+1}}={\pazocal{O}}\left(\frac{1}{\sqrt{k+1}}\right),$ (3.27) where $\kappa_{\mbox{\tiny ADSW}}=\kappa_{T}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right)+\frac{4\kappa_{N}\,\kappa_{g}}{\beta\eta}.$
   • Proof.
+Observe first that AS.8, AS.9 and AS.10 automatically hold in the gradient is exact.
 Now consider iterations from 0 to $k$ of both types (tangential and normal) by setting $\min[k_{\nu_{0}},k_{\tau_{0}}]=0$ and $\max[k_{\nu_{1}},k_{\tau_{1}}]=k$ (as in Lemma 3.5).
 We then obtain, by combining (3.20) and (3.22), that
-    $\sum_{j=0}^{k}$ $\big(\|g_{T,j}\|+\|c_{j}\|\Big)=\sum_{\tau=\tau_{0}}^{\tau_{1}}\Big(\|g_{T,k_{\tau}}\|+\|c_{k_{\tau}}\|\Big)+\sum_{\nu=\nu_{0}}^{\nu_{1}}\Big(\|g_{T,k_{\nu}}\|+\|c_{k_{\nu}}\|\Big)$
-    $\leq\kappa_{T}\sqrt{k+1}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right)+\kappa_{N}\,\left(1+\frac{2\sqrt{2}}{\beta\eta}\right),$
-    where we used the inequalities $\tau_{1}\leq k_{\tau_{1}}\leq k$ and $k_{\nu_{1}}\leq k$. The bound (3.26) is finally obtained by dividing both sides by $k+1$. $\Box$
+    $\sum_{j=0}^{k}$ $\mathbb{E}\!\left[\|G_{T,j}\|+\|c_{j}\|\right]=\sum_{\tau=\tau_{0}}^{\tau_{1}}\mathbb{E}\!\left[\|G_{T,k_{\tau}}\|+\|c_{k_{\tau}}\|\right]+\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|G_{T,k_{\nu}}\|+\|c_{k_{\nu}}\|\right]$
+    $\leq\kappa_{T}\sqrt{k+1}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right)+\frac{4\kappa_{N}\,\kappa_{g}}{\beta\eta}\sqrt{k+1},$
+    where we used the inequalties $\tau_{1}\leq k_{\tau_{1}}\leq k$ and $k_{\nu_{1}}\leq k$. The bound (3.27) is finally obtained by dividing both sides by $k+1$. $\Box$
+We may now use a standard probabilistic argument and show that the probability of a small optimality measure $\|G_{T,k}\|+\|c_{k}\|$ increases to one when $k$ grows.
+Corollary 1
+Suppose that AS.0–AS.10 hold.
+Then $\mathbb{P}\!\left[\min_{j\in\{0,\ldots,k\}}\left[\|G_{T,j}\|+\|c_{j}\|\right]\leq\epsilon\right]\geq 1-\delta\;\;\mbox{for}\;\;k>\frac{\kappa_{\mbox{\tiny ADSW}}^{2}}{\epsilon^{2}\delta^{2}}.$
+  • Proof.    Markov’s inequality and (3.27) give that
+    $\mathbb{P}\!\left[\min_{j\in\{0,\ldots,k\}}\Big(\|G_{T,j}\|+\|c_{j}\|\Big)\leq\epsilon\right]$ $\geq\mathbb{P}\!\left[\frac{1}{k+1}\sum_{j=0}^{k}\Big(\|G_{T,j}\|+\|c_{j}\|\Big)\leq\epsilon\right]$
+    $\geq 1-\frac{1}{\epsilon}\mathbb{E}\!\left[\sum_{j=0}^{k}\Big(\|g_{T,j}\|+\|c_{j}\|\Big)\right]$
+    $\geq 1-\frac{\kappa_{\mbox{\tiny ADSW}}}{\epsilon\sqrt{k+1}}.$
+    The desired conclusion follows. $\Box$
 4 Numerical illustration
@@
     The algorithm appears to converge as predicted.
-Despite the fact that our current analysis does not cover the case where the Jacobian may lose rank, the algorithm did solve a number of instances where this occurs, sometimes finding an infeasible critical point of the constraints’ violation [^5] This occurs, for instance, at the starting point of HS61..
+Despite the fact that our current analysis does not cover the case where the Jacobian may loose rank, the algorithm did solve a number of instances where this occurs, sometimes finding an infeasible critical point of the constraints’s violation [^5] This occurs, for instance, at the starting point of HS61..
   2. 2.
@@
 This is also the case for ADSWITCH.
-Over the 71 problems in our test set, it solves 44 (62%) of them within 750 iterations, 58 (81%) of them within 100000 iterations and fails on 13.
+Over the 71 problems in our test set, it solves 44 of them within 750 iterations, 58 of them within 100000 iterations and fails on 13.
 While reliability is not a strong point of AdaGrad and ADSWITCH in the deterministic case, the situation is very different when noise is present, as we show below.
 We illustrate a few case of satisfactory convergence in Figures 1 and 2.
-In these figures, one can clearly see the difference in speed of convergence between tangential steps (AdaGrad-like) and normal steps (Newton), and thus that the overall performance is dominated by that of the first-order method defining the tangential step.
+In these figures, one can clearly see the difference in speed of convergence between tangential steps (AdaGrad-like) and normal steps (Newton), and thus that the overal performance is dominated by that of the first-order method defining the tangential step.
 This is even clearer when convergence is too slow, as shown in Figure 4 where constraint violation remains very small [^6] HS50 has linear equality constraints only., while very slow convergence of the projected gradient results in a large number of iterations.
-The left panel of Figure 1 (ORTHREGA with $n=133$ and $m=64$) is also interesting because it shows the switching condition (2.12) in action: the algorithm first approaches an infeasible critical point of the constraints’ violation until the switching rule allows the tangential step to take over, causing the iterates to escape.
+The left panel of Figure 1 (ORTHREGA with $n=133$ and $m=64$) is also interesting because it shows the switching condition (2.12) in action: the algorithm first approaches an infeasible critical point of the constraints’s violation until the switching rule allows the tangential step to take over, causing the iterates to escape.
 Figure 1: Evolution of the optimality measure (left: BYRDSPHR, right: ORTHREGA)
@@
 This stable behaviour is also visible in Figure 5 showing the performance of the algorithm on the BT1 problem for increasing noise levels.
-Figure 5: Evolution of the two components of the optimality measure for a run of the noisy BT1 problem, with 5% (top left), 15% (top right), 25% (bottom left) and 50% (bottom right) relative Gaussian noise on the gradient
+Figure 5: Evolution of the two components of the optimality measure for a run of the the noisy BT1 problem, with 5% (top left), 15% (top right), 25% (bottom left) and 50% (bottom right) relative Gaussian noise on the gradient
 5 Conclusions and perspectives
-false We have proposed a very simple first-order algorithm for solving nonlinear optimization problems with nonlinear equality constraints.
+We have proposed a very simple first-order algorithm for solving potentially stochastic nonlinear optimization problems with deterministic nonlinear equality constraints.
 This algorithm adaptively selects steps in the plane tangent to the constraints or steps that reduce infeasibility.
@@
 As is the case for AdaGrad, the objective function is never evaluated.
-We have analyzed its worst-case evaluation complexity and obtain global convergence rates which match the best known rates for unconstrained problems.
+We have analyzed its worst-case evaluation complexity, both in the deterministic and stochastic cases, and obtain global convergence rates which match the best known rates for unconstrained problems.
 Numerical experiments have been presented indicating that the performance of the algorithm is comparable to that of first-order methods for unconstrained problems.
 At this stage, several theoretical questions remain open for further research.
-These include a stochastic convergence analysis, covering the cases of rank-deficient Jacobians and/or unbounded gradients, the use of alternative minimization methods in the tangential step (such as ADAM [23], ASTR1 [19], OFFAR [20] or more standard unconstrained minimizing techniques using objective function’s values).
+These include covering the cases of rank-deficient Jacobians and/or unbounded gradients, the use of alternative minimization methods in the tangential step (such as ADAM [21], ASTR1 [17], OFFAR [18] or more standard unconstrained minimizing techniques using objective function’s values).
 On the practical side, further numerical experiments are clearly necessary for assessing the true potential of the new method.
@@
 Acknowledgement
-Philippe Toint is grateful for the continued and friendly support of the Algo team at Toulouse IRIT. The authors also thank Benedetta Morini and Stefania Bellavia for pointing out a number of typos.
+Philippe Toint is grateful for the continued and friendly support of the Algo team at Toulouse IRIT.
+Both authors thank Mehi Al-Baali and the organizers of the NAOVI 2026 conference for their kind invitation to participate.
 Appendix; Details of the Numerical Results
@@
 Columns 2 and 3 indicate the number of variables and the number of constraints.
-The column “#its” gives the number of iterations and the column “exitc” the termination condition (“convg” = convergence, “infeas” = infeasible critical point, “maxit” = maximum number of iterations reached).
+The colmun “#its” gives the number of iterations and the column “exitc” the termination condition (“convg” = convergence, “infeas” = infeasible critical point, “maxit” = maximum number of iterations reached).
   • Tables A.2, A.3, A.4 and A.5 report the average statistics of 10 independent runs for each problem, for increasing relative Gaussian noise levels.
@@
 SPINOP 11 9 +8.435060e-02 7.91e-04 2.20e-06 100000 maxit
-Table A.1: Results of running ADSWITCH in the abscence of noise
+Table A.1: Results of running ADAGEC in the abscence of noise
 Problem $n$ $m$ avr $f(x)$ avr $\|g_{T}(x)\|$ avr $\|c(x)\|$ avr #its #success
@@
 SPINOP 11 9 +9.941025e-02 1.10e-03 5.74e-04 6.14e+04 10
-Table A.2: Results of running ADSWITCH with 5% relative Gaussian noise
+Table A.2: Results of running ADAGEC with 5% relative Gaussian noise
 Problem $n$ $m$ avr $f(x)$ avr $\|g_{T}(x)\|$ avr $\|c(x)\|$ avr #its #success
@@
 SPINOP 11 9 +1.134528e-01 1.43e-03 6.21e-04 4.12e+04 10
-Table A.3: Results of running ADSWITCH with 15% relative Gaussian noise
+Table A.3: Results of running ADAGEC with 15% relative Gaussian noise
 Problem $n$ $m$ avr $f(x)$ avr $\|g_{T}(x)\|$ avr $\|c(x)\|$ avr #its #success
@@
 SPINOP 11 9 +1.331340e-01 1.97e-03 5.60e-04 2.56e+04 10
-Table A.4: Results of running ADSWITCH with 25% relative Gaussian noise
+Table A.4: Results of running ADAGEC with 25% relative Gaussian noise
 Problem $n$ $m$ avr $f(x)$ avr $\|g_{T}(x)\|$ avr $\|c(x)\|$ avr #its #success
@@
 SPINOP 11 9 +3.658931e-01 1.59e-02 4.10e-04 1.53e+03 10
-Table A.5: Results of running ADSWITCH with 50% relative Gaussian noise
+Table A.5: Results of running ADAGEC with 50% relative Gaussian noise